### v1

# DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric Calibration

Abstract
We use hierarchical Bayesian modelling to calibrate a network of 32 all-sky faint DA white dwarf (DA WD) spectrophotometric standards ($19.516.5<V<$ 19.5) alongside the three CALSPEC standards, from 912 Å to 32 $\mu$m. The framework is the first of its kind to jointly infer photometric zeropoints and WD parameters ($\log g$, $T_{\text{eff}}$, $A_{V}$, $R_{V}$) by simultaneously modelling both photometric and spectroscopic data. We model panchromatic HST/WFC3 UVIS and IR fluxes, HST/STIS UV spectroscopy and ground-based optical spectroscopy to sub-percent precision. Photometric residuals for the sample are the lowest yet yielding $<0.004$ mag RMS on average from the UV to the NIR, achieved by jointly inferring time-dependent changes in system sensitivity and WFC3/IR count-rate nonlinearity. Our GPU-accelerated implementation enables efficient sampling via Hamiltonian Monte Carlo, critical for exploring the high-dimensional posterior space. The hierarchical nature of the model enables population analysis of intrinsic WD and dust parameters. Inferred SEDs from this model will be essential for calibrating the James Webb Space Telescope as well as next-generation surveys, including Vera Rubin Observatory’s Legacy Survey of Space and Time, and the Nancy Grace Roman Space Telescope.
keywords:
methods: statistical – white dwarfs – standards – surveys – stars: statistics – stars: fundamental parameters
pubyear: 2024pagerange: DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric CalibrationI
1 Introduction
As we approach a new era of precision cosmology, our systematic uncertainties will outweigh our statistical uncertainties. This effect will be further amplified with the next-generation of wide-field surveys, such as the Legacy Survey of Space and Time at the Vera Rubin Observatory (LSST; Ivezić et al., 2019), as Type Ia supernova (SN Ia) sample sizes increase to over a million.
A particularly delicate systematic uncertainty to consider is the spectrophotometric calibration of surveys (Stubbs & Brown, 2015). Calibration can be absolute, aiming to define the flux of an object in relation to International System of Units (Bohlin, 2014; Gordon et al., 2022). Alternatively, the calibration can be relative, where an arbitrary flux scale is chosen, but the objective is to characterise an object’s change in flux as a function of wavelength (Narayan et al., 2019; Axelrod et al., 2023). In this work, we introduce DAmodel: a hierarchical Bayesian framework for accurately calibrating the relative flux of 32 faint DA white dwarfs, along with three primary standards. The model utilises photometric and spectroscopic data from diverse sources, achieving unprecedented levels of flux calibration, ready for the demands of next-generation surveys.
There is a particularly strong degeneracy between the photometric redshift of a galaxy and the relative calibration of a survey’s photometric zeropoints (High et al., 2009; Euclid Collaboration et al., 2024). If a survey’s relative zeropoints are calibrated incorrectly, the difference in flux could be misinterpreted as additional colour that would then compromise the photometric redshift estimate. This effect can also influence the Hubble diagram in SN Ia cosmology. Although systematic effects only change colours of individual objects by the order of a percent (Betoule et al., 2013), the combined effect of incorrectly measuring an entire SN population’s colour would have a significant influence on cosmological constraints (Hounsell et al., 2018). Scolnic et al. (2014) showed that colour calibration systematics can create an additional signal on the Hubble diagram at the 0.01-0.015 magnitude level. In the DESI analysis (DESI Collaboration et al., 2024), the difference between an evolving $w(z)$ and constant $w=-1$ dark energy equation of state Universe was a signal of only 0.04 mag, emphasising the importance of having confidence in relative photometric calibration. A further process in supernova cosmology that requires accurate relative flux calibration is the comparison of low redshift and high redshift samples. These contrasting samples are necessary to measure the change in the relative curve of the Hubble diagram, however, if these samples use inconsistent definitions of brightness or colour then they will be at odds with our true underlying cosmology (Efstathiou, 2024; Dhawan et al., 2024). To overcome this issue, it is imperative to perform reliable cross-calibration between surveys that rely on accurate flux standards.
Figure 1: Coordinates and Gaia $G$ band brightness of the DA white dwarfs in our network. The red stars illustrate the CALSPEC primary standards that have brightnesses 11.7< $G$ <13.3.
Spectrophotometric calibration relies on the precise modelling of the spectral energy distributions (SEDs) of astronomical objects that are well understood. Pure-hydrogen atmosphere DA white dwarfs (DA WDs) are ideal candidates to be modelled as standards, since they are near-perfect black bodies, with the addition of strong absorption features that are correlated with surface gravity and effective temperature (Holberg et al., 1985; Holberg & Bergeron, 2006; Hubeny & Lanz, 2017). So far, the three most accurate DA WDs used for spectrophotometric calibration have been GD71, GD153 and G191B2B. These WDs have little reddening and have been staples of the CALSPEC system (Bohlin et al., 2014, 2020), allowing for the calibration of astronomy’s most successful space missions, such as the Hubble Space Telescope (HST) and Gaia (Gaia Collaboration et al., 2016), as well as various other ground-based instruments. Since different missions scan different regions of the sky, it is necessary for networks of standards to have coverage over large portions of the celestial sphere. Rubin et al. (2022a) was awarded 35 HST/STIS orbits to observe 11 DA WDs to be added to CALSPEC system. Elms et al. (2024) recently expanded the CALSPEC standards by adding 17 new cool ($T_{\text{eff}}$ < 20,000 K) DA WDs. Low temperature DA WDs are particularly useful for standardisation as they have reduced non-local thermal equilibrium effects (more predictable SEDs) and are brighter in the near-infrared (NIR).
An essential factor in spectrophotometric calibration that must be treated with care is the modelling of the effects of interstellar dust within the Milky Way. Dust along the line of sight between an object and the Earth has the combined effect of extinguishing and reddening light (Draine, 2003). Extinction, often measured in the $V$ band as $A_{V}$, is dependent on the column density of dust that lies between object and observer (e.g. Foster et al., 2013), which is also correlated with distance. The quantity $R_{V}=A_{V}/E(B-V)$, where $E(B-V)$ is the colour excess between $B$ band and $V$ band, is used to determine the amount of reddening one would expect from different types of dust grains. There have been various functional forms of dust relations that use these parameters to define extinction as a continuous function of wavelength $A(\lambda;A_{V},R_{V})$ (e.g. Cardelli et al., 1989; Fitzpatrick, 1999; Fitzpatrick et al., 2019; Gordon et al., 2023), that have been a key ingredient in spectrophotometric modelling DA WDs. Since DA WDs are so well understood, they are also useful tools to study dust. Sahu et al. (2024) recently used DA WDs to construct three-dimensional dust maps in the UV.
The value of $A_{V}$ has a strong impact in the UV wavelengths, particularly around the 2175 Å bump that leads to a sharp drop off in extinction (Fitzpatrick & Massa, 1986). For this reason, it is important to spectroscopically observe DA WDs in the UV to constrain $A_{V}$ and extrapolate SEDs with a higher degree of confidence to longer wavelengths in the thermal and far infrared. Dust extinction decreases significantly in the NIR, meaning that if we have data coverage at these wavelengths we can disentangle degenerate dust and intrinsic parameters. The difference in extinction at UV and NIR wavelengths allows for more reliable inference of the $R_{V}$ parameters. This panchromatic strategy to mitigate dust effects is also used elsewhere in astronomy, for example in SN Ia modelling (Mandel et al., 2022; Thorp & Mandel, 2022; Grayling et al., 2024; Thorp et al., 2024a; Grayling & Popovic, 2024) and inference of galaxy properties Alsing et al. (2024); Thorp et al. (2024b).
As we look towards the next-generation of deep surveys such as LSST (Ivezić et al., 2019), the Nancy Grace Roman Space Telescope (Roman; Green et al., 2012) and the James Webb Space Telescope (JWST; Gardner et al., 2006), we need dimmer standards to calibrate them. The three primary CALSPEC DA WDs, with magnitudes ranging from 11.7 < $V$ < 13.4, have been observed with JWST but are close to its observational limit and exceed the normal observing mode saturation limits of upcoming surveys, motivating the need for a new set of fainter standards. Our collaboration (led by Co-PIs Abhijit Saha and Gautham Narayan) has established an all-sky network of 32 faint (16.5 < $V$ < 19.5) DA white dwarfs to be used as spectrophotometric standards (Narayan et al., 2016, 2019; Calamida et al., 2019, 2022b; Axelrod et al., 2023). This group of DA WDs, illustrated in Figure 1, is intended to deliver exceptionally precise flux calibration, ready for the upcoming era of deep-sky surveys.
Narayan et al. (2019) developed a novel Bayesian methodology to model the faint WDs using HST photometry paired with ground-based optical spectroscopy (Calamida et al., 2019) that would go on to achieve sub-percent accuracy. The approach consisted of a two-step process. First, HST/WFC3 photometry was calibrated for each WD across six bands, spanning from the UV to the NIR. This step included addressing various instrumental systematics, such as time-dependent changes in zeropoint sensitivity in each band (Calamida et al., 2022a; Marinelli et al., 2024). The second step of the process took the calibrated photometry and ground based optical spectra to serially model SEDs of each WD by inferring intrinsic and dust parameters.
The work of Narayan et al. (2019) was particularly successful in the optical and UV wavelengths, yielding an average RMS of 4 mmag across the five bands. The NIR F160W band proved to be more difficult due to known count-rate nonlinearity (CRNL) systematics (Riess et al., 2019; Bohlin & Deustua, 2019). Axelrod et al. (2023) subsequently expanded and reanalysed the network by using the approach developed by Narayan et al. (2019) to establish spectrophotometric priors on the model parameters. They then transitioned to a hierarchical Bayesian model to address the CRNL more accurately than using photometry alone. This combination proved successful, accomplishing a RMS of 8 mmag in the NIR, demonstrating the benefit of hierarchically analysing the network. The hierarchical approach enabled the inference of a band-dependent zeropoint shift relative to the CALSPEC system, which can only be derived from a set of stars. This allowed relative independence from CALSPEC which was crucial for ensuring the network could be calibrated to any magnitude system. It became clear that the next objective for this network would be a fully hierarchical end-to-end Bayesian analysis that would combine all spectroscopic and photometric data available.
Hierarchical Bayesian models (HBMs) have become a powerful tool in astronomy, offering a flexible way to handle complex functional data with multiple layers of uncertainty (e.g. Loredo et al., 2024). Their strength lies in the ability to model dependencies across different levels, such as population parameters for groups of objects and individual-level parameters that vary within these groups. This hierarchical approach enables astronomers to account for intrinsic scatter, observational uncertainties, and selection biases, which are common in astronomical surveys. HBMs have been applied across many areas of astronomy for a variety of applications, including exoplanet atmospheric retrievals (Lustig-Yaeger et al., 2022), stellar evolution (Si et al., 2017, 2018), galaxies (Loredo & Hendry, 2019; Leistedt et al., 2023), gravitational wave analysis (Thrane & Talbot, 2019), SN Ia modelling (Mandel et al., 2009, 2011, 2014, 2017, 2022) and cosmology (March et al., 2011; Rubin et al., 2015; Feeney et al., 2018; Boyd et al., 2024). A HBM was used in Rubin et al. (2022b) to perform a simultaneous spectrophotometric recalibration of 32 stars, including 14 existing CALSPEC standards, based on measurements obtained with the SuperNova Integral Field Spectrograph. The work was able to achieve 14 mmag accuracy by carefully modelling population-level instrumental systematics relating to atmospheric extinction, temperature dispersion and nightly calibration. In addition, HBMs were also used for the calibration of X-ray surveys in Chen et al. (2019) and Marshall et al. (2021).
In this work we establish DAmodel: a scalable fully hierarchical Bayesian framework for spectrophotometric calibration of our all-sky DA WD network (including 32 faint and 3 CALSPEC standards) to the highest accuracy yet, allowing for precision astronomy in the next-generation of surveys. For the first time, we jointly analyse HST/WFC photometry in five UVIS broad bands and one NIR band, high resolution optical ground-based spectra (Calamida et al., 2019, 2022b) and new UV spectra from the Hubble Space Telescope Imaging Spectrograph (STIS). This work solves the calibration problem by implementing a novel multi-level hierarchy that can use the network to simultaneously infer instrumental systematics, dust relations and intrinsic WD parameters.
In Section 2, we outline the spectrophotometric DA WD dataset used in this work. We then outline the statistical model of the DAmodel framework in Section 3. In Section 4 we present our calibrated SEDs, synthetic photometric residuals with HST and other surveys, then discuss the implications of our results. Future applications of the framework are discussed in Section 5 and we conclude with final remarks in Section 6.
Figure 2: Visualisation of available data for WDFS0458-56. The coloured circles represent HST/WFC3 UVIS and IR photometry, increasing in wavelength from left to right. Highlighted in green is the ground-based optical spectrum from the Southern Astrophysical Research (SOAR) telescope. Shaded in pink is the new HST/STIS UV spectrum that we have also obtained for 18 other faint DA WD sources. Under-plotted in black is our SED model evaluated at the parameters inferred for this particular source.
2 Data
In Section 2.1 we describe how the standards were selected. In Section 2.2 and Section 2.3 we briefly describe the data previously analysed in Narayan et al. (2019) and Axelrod et al. (2023). More in-depth detail on the dataset can be found in Calamida et al. (2019) and Calamida et al. (2022b). Finally, in 2.4 we outline the data collection and reduction used for the new HST/STIS UV spectra.
2.1 Target Selection
Probable DA WDs in the northern and equatorial regions were selected by Narayan et al. (2019) and Calamida et al. (2019) using known sources, including Sloan Digital Sky Survey (SDSS) observations (Kleinman et al., 2004; Eisenstein et al., 2006) and the earlier Villanova catalogue (McCook & Sion, 1999), which has since been updated to the Montreal database (Dufour et al., 2017). These DA WD candidates had previously published low-resolution spectra, which provided enough detail to confirm them as DA type, assess their radiative properties, and check for any confounding factors such as magnetic line splitting or trace atmospheric elements. In contrast, no comprehensive list existed for faint DA WDs in the southern hemisphere, so candidates were identified from the SuperCOSMOS and VST surveys (Raddi et al., 2016, 2017) using photometric data and proper motion selection, focusing on stars with absolute magnitudes brighter than $V=$9.0. Only DA WDs with effective temperatures above 19,500 K were kept to be within the bounds where we have confidence in theoretical templates. It was also important to cut sources that are strongly influenced by magnetic fields, as these effects are not included in our modelling. This influence will present itself via Zeeman splitting of the Balmer lines, or in the presence of the Minkowski bands, shallow and broad absorption bands near 3650 Å, 4135 Å, and 4466 Å. Using the optical spectra, DA WDs were cut that exhibited strong magnetic traits, however, those with weaker effects could not be thrown away due to degeneracy with the surface gravity parameter.
The main goals of the study described in Calamida et al. (2022b) were to display spectra for all stars observed with the Southern Astrophysical Research (SOAR) telescope and to examine these stars for variability using the Las Cumbres Observatory (LCO) telescope network. Hot DA white dwarfs are not expected to vary intrinsically, but variability could arise due to binary companions, “seeing” effects from nearby faint red stars, or dust clouds surrounding the white dwarf. Narayan et al. (2019) and Calamida et al. (2019) eliminated four candidate stars based on spectroscopic and photometric data, while Calamida et al. (2022b) excluded a total of six stars from the full-sky set, leaving 32 faint stars and three brighter CALSPEC standards to establish our network. Figure 1 illustrates the locations and brightness of our faint all-sky network.
2.2 Hubble Space Telescope WFC3 Photometry
The optical photometry used in this study is extensively documented in Calamida et al. (2019) and Calamida et al. (2022b), while here we summarize the primary details of its collection. The photometry was acquired using the WFC3 with the UVIS and IR detectors, as part of Cycle 20 (average MJD=56452.94 days) and Cycle 22 (average MJD=57076.11 days), led by PI Abhijit Saha (proposals GO-12967 and GO-13711). Cycle 20 observations utilized five filters: F336W, F475W, F625W, F775W, and F160W. In Cycle 22, an additional near-UV (NUV) filter, F275W, was included to enhance detection of line-of-sight extinction. Some DA WD samples were only observed in one cycle, while others were observed in both. The latter group included nine DA WDs near the celestial equator. Observing this sample in both cycles not only improved photometric precision but also allowed for comparison between the two observing epochs.
Cycle 20 observations of the equatorial DA WDs were also used to validate the methodology of establishing spectrophotometric standards based on earlier-collected DA WDs (Narayan et al., 2016). Additionally, 14 candidate DA WDs and three CALSPEC WDs were observed during Cycle 22 (Bohlin et al., 2014, 2020), tying all observations to the HST photometric system. Collection methods in Cycle 20 varied by filter: F336W and F475W had three dithered exposures, while the other three filters had two exposures each. In Cycle 22, equatorial DA WDs were observed with three F275W exposures (with cosmic ray splits) and one additional exposure in each Cycle 20 filter. Ensuring at least three exposures per band for each source allowed for cross-exposure comparison, addressing issues like cosmic rays and detector artifacts in HST WFC3 images and maintaining sub-percent photometric precision. For WDs observed only in Cycle 22, five to seven dithered exposures per filter were obtained. The CALSPEC WDs were observed on three separate visits, with six to eight exposures per filter per visit. Observations in Cycle 20 used the UVIS1 sub-arrays and remained on the UVIS1 photometric system, while Cycle 22 observations used UVIS2 sub-arrays and were on UVIS2 photometric system.
The southern DA white dwarfs were observed with three HST Cycle 25 (average MJD=57588.79 days) orbits per star using WFC3 (Calamida et al., 2022b). The three main CALSPEC spectrophotometric standards GD71, GD153, and G191B2B each had three visits in Cycle 25 to address small potential sensitivity variations of WFC3. In 1.5 years, we expect only minor sensitivity variations, as sensitivity loss rates range from 0.1% to 0.2% per year, with greater losses at bluer wavelengths (Calamida et al., 2022a). Final photometric reductions were performed using Saha’s ILAPH code, as detailed in Calamida et al. (2019). This custom aperture photometry tool includes an interactive "growth curve" analysis for optimal sky subtraction. Ensuring the best possible count rates from the images was crucial, but equally important was ensuring that all measurements were placed on a consistent system of instrumental magnitudes. During post-processing, Cycle 25 observations were converted into the UVIS1 photometric system in the calwf3 calibration pipeline. Given that the HST/WFC3 data were obtained at different epochs over several years and with varying instrument configurations, special attention was given to adjusting for potential systematic magnitude shifts. The final photometry was placed on the AB magnitude system of CALSPEC. Although the use of AB magnitudes is not critical for subsequent analysis, it provided a convenient and self-consistent reference system. Ultimately, this process led to the determination of new AB magnitudes that are independent of the initial measurement system used for the observations.
2.3 Ground-Based Optical Spectra
Ground-based spectra were obtained with the Gemini North and South telescopes, the Multiple Mirror Telescope (MMT) and SOAR, as described in Calamida et al. (2019) and Calamida et al. (2022b). Spectroscopic data were obtained using the Gemini Multi-Object Spectrograph (GMOS) at both Gemini North and South observatories (Calamida et al., 2019). Observations were conducted in queue mode, initially with a $1.5^{\prime\prime}$ slit, which was later narrowed to a $1.0^{\prime\prime}$ slit. The final Gemini spectra cover a continuous range from 3500 Å to 6360 Å with a dispersion of 0.92 Å/pixel (Calamida et al., 2019). However, it was determined that these spectra were insufficient for use as spectrophotometric standards due to several factors: instrumental limitations (such as GMOS’s limited blue throughput), the absence of standard CALSPEC stars (Bohlin et al., 2014, 2020) in the observing runs, and a lack of parallactic angle alignment during observations, resulting in atmospheric dispersion losses (Calamida et al., 2019). This led to further observations at the MMT Observatory to enhance the spectral data.
At MMT, DA WDs were observed using the Blue Channel spectrograph with a 300 line $\text{mm}^{-1}$ grating. Most observations employed a $1.0^{\prime\prime}$ slit, with one epoch using a $1.25^{\prime\prime}$ slit (Calamida et al., 2019). The MMT spectra cover a wavelength range from 3400 Å to 8400 Å at a dispersion of 1.95 Å/pixel. All observations at MMT were taken at the parallactic angle, and standard stars were observed on the same night. The spectrographs did not operate on a linear pixel-to-angstrom scale, so the data were resampled onto a uniform linear grid after extraction, with a sampling of 1 Å/pixel for GMOS and 2 Å/pixel for MMT (Calamida et al., 2019). Finally, the spectra were flux-calibrated using custom routines designed for calibrating SNe Ia spectra (Matheson et al., 2008). Flux calibration was performed using a low-order polynomial with knot spacing of several hundred angstroms, making it insensitive to the sampling scale of 1–2 Å.
Approximately 50 southern white dwarf candidates were observed using the Goodman spectrograph at SOAR ($1.07^{\prime\prime}$ slit, $1.99^{\prime\prime}$/pixel, 3850–7100 Å coverage) over several runs in 2016 and 2017. A list of candidates observed and a log of observations are in (Calamida et al., 2022b).
2.4 Space-Based STIS UV Spectra
To better characterize and ultimately correct for dust effects, the spectroscopic observations were extended into the UV regime. This expansion was critical for capturing the 2175 Å bump in the dust relation, where dust extinction is most pronounced. These UV observations aimed to identify potential "anomalous" reddening sources that deviate from current extinction relations. To accurately quantify extinction and minimize observational noise, the HST Space Telescope Imaging Spectrograph (STIS) was used for data collection. The target selection criteria included the brighter standards, with magnitudes of 15 $\leq$ F275W $\leq$ 18, and known $A_{V}$ values between 0 and 1.
Spectra for 21 WDs were obtained using the STIS FUV and NUV detectors with the G140L and G230L gratings, respectively, under program 16764 led by PI Gautham Narayan. Both FUV and NUV observations employed a $52\times 2$ arcsecond slit width, covering wavelength ranges of 1150 to 1700 Å for FUV and 1570 to 3180 Å for NUV. FUV exposure times varied from 732 to 2518 seconds, while NUV exposure times ranged from 70 to 2788 seconds. Each orbit’s visit structure included an acquisition (ACQ) image, with an acquisition time between 4 seconds and 40 seconds. One orbit was used for the brightest WD, while 2 to 3 orbits were required for fainter, higher-extinction sources.
After the variability selection described in Section 2.1, 19 out of our 32 faint standards have STIS UV spectra to complement their optical ground-based spectra. The specific WDs that had this additional coverage are indicated in Table 1 and Table 3. Further information on the STIS UV spectra can be found in the preliminary analysis by Bohlin et al. (2024).
Figure 3: Probabilistic graphical model (directed acyclic graph) illustrating the relationships between parameters in our hierarchical Bayesian model. In the diagram we see multiple levels of of hierarchy where the top level population dust parameters ($\mu_{RV}$,$\sigma_{RV}$,$\tau$) influence each WD’s extinction. The object-level parameters ($\mu_{s}$,$T_{\text{eff}}^{s}$, $\log g^{s}$,$A_{V}^{s}$,$R_{V}^{s}$) determine the SED of each individual WD. The nuisance parameters ($\bm{\epsilon}^{j}_{s}$,$\text{FWHM}^{j}_{s}$,$\gamma^{j}_{s}$) allow us to model spectra $\bm{\hat{S}}^{j}_{s}$ from the each object’s SED. There are systematic parameters ($Z_{X}$,$\alpha_{X}$,$\sigma^{X}_{\text{int}}$,$\nu^{X}$) in each band $X$, as well as HST cycle-dependent offsets $\Delta Z^{c}_{X}$, allowing us to model photometric observations $\hat{m}_{i,c,s}$ of the network. Each parameter is defined clearly in Section 3.
3 The Statistical Model
In this section we define the DAmodel statistical model, building on techniques previously developed by Narayan et al. (2019) and Axelrod et al. (2023). In Section 3.1 we describe the forward model to construct our dusty SEDs. In Section 3.2 we describe how we use the dusty SEDs in our likelihoods to model the HST/WFC photometry and instrumental systematics. In Section 3.3 we describe how we adapt the dusty SEDs to model observed optical and STIS UV spectra. In Section 3.4 we describe how we combine the different elements of the statistical model to make a fully hierarchical Bayesian model, which is also illustrated in Figure 3. Finally, in Section 3.5 we define the choice of non-informative priors on each inferred parameter.
3.1 Synthetic Spectral Energy Distribution
The model requires an underlying grid to describe the relationship between the spectral energy distribution (SED) of the DA white dwarf with intrinsic parameters describing effective temperature $T_{\text{eff}}$ and the natural logarithm of the surface gravity $\log g$. The distinct depths of the Balmer series in the optical wavelengths are strongly correlated with $T_{\text{eff}}$, whereas their equivalent widths are dependent on $\log g$ (Tremblay & Bergeron, 2009).
[…]
The Hubeny grid is defined from 900 Å to 320,000 Å with resolution of $R=$100,000 for DA WDs in the range of 15,000 K < $T_{\text{eff}}$ < 70,000 K and 7 < $\log\text{g}$ < 9.5. Our model’s unreddened SED $\mathcal{F}(\lambda;\log g,T_{\text{eff}})$ is determined by the cubic interpolation of the theoretical grid. The SED is then reddened according to the Gordon et al. (2023) dust relation where the extinction $A(\lambda;A_{V},R_{V})$ is applied such that
$\tilde{F}(\lambda;\log g,T_{\text{eff}},A_{V},R_{V})=\mathcal{F}(\lambda;\log g,T_{\text{eff}})\cdot 10^{-0.4\cdot{A(\lambda;A_{V},R_{V}})}$ (1)
where $A_{V}$ is the extinction in $V$ band and $R_{V}=A_{V}/E(B-V)$, where $E(B-V)$ is the colour excess between $B$ band and $V$ band. The extinction relation is defined between 912 Å and 32 $\mu$m which is what dictates the limits of our published SEDs. In Section 4.5 we discuss how using Gordon et al. (2023) dust relation over Fitzpatrick (1999), that was used in previous analyses, affects our results.
We apply an achromatic offset to the reddened SED using
$F(\lambda;\log g,T_{\text{eff}},A_{V},R_{V},\mu)=\tilde{F}(\lambda;\log g,T_{\text{eff}},A_{V},R_{V})\cdot 10^{-0.4\mu}$ (2)
The pseudo-distance modulus $\mu$ is a scale factor expressed in magnitudes with an arbitrarily defined zeropoint, relating an object’s surface brightness to its apparent brightness. The value of $\mu$ depends on both the physical size of the object and its distance from the observer. In our analysis methodology, neither the physical size nor the distance is determined individually, as we do not impose a mass-radius-luminosity relation to derive the absolute brightness.
3.2 Photometric Modelling
[…]
where $F_{\text{AB}}(\lambda)$ is the spectral flux density of the fiducial AB source expressed in $\text{ergs cm}^{-2}$ $\text{s}^{-1}$$\text{\AA}^{-1}$. Intrinsic WD and dust parameters collectively represented by $\bm{\Theta^{s}_{\text{WD}}}=(\log g^{s},T^{s}_{\text{eff}},A^{s}_{V},R^{s}_{V})^{T}$.
We then construct a likelihood to model individual photometric observations $\hat{m}^{X}_{i,c,s}$ of WD $s$ in band $X$ from Hubble Cycle $c$ as
$P\big{(}\hat{m}^{X}_{i,c,s}|\ \bm{\Theta^{s}_{\text{WD}}},\mu_{s},\bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}=\\ T\big{(}\hat{m}^{X}_{i,c,s}|\ m^{X}_{s}-Z_{X}-\Delta Z_{X}^{c}-\alpha_{X}(m^{X}_{s}-15),\sqrt{\sigma^{X\hskip 8.5359pt2}_{s,i,c}+\sigma_{\text{int}}^{X\hskip 2.84544pt2}},\nu^{X}\big{)}$ (4)
where $T(x|\ \bar{\mu},\sigma,\nu)$ is a student-T distribution with mean $\bar{\mu}$, scale $\sigma$ and $\nu$ degrees of freedom. We choose a student-T distribution to model the photometry due to its broader tails, allowing flexibility and robustness in modelling outliers. The measurement error for each individual photometric observation $i$ is represented by $\sigma^{X}_{s,i,c}$. Each HST band $X$ has its own set of parameters $\bm{\theta_{\text{phot}}^{X}}=(Z_{X},\sigma_{\text{int}}^{X},\nu^{X},\alpha_{X})^{T}$ containing a zeropoint $Z_{X}$, excess dispersion $\sigma_{\text{int}}^{X}$, degree of freedom $\nu^{X}$ and CRNL coefficient $\alpha_{X}$ (Riess et al., 2019; Bohlin & Deustua, 2019; Axelrod et al., 2023). The excess dispersion parameter $\sigma_{\text{int}}^{X}$, first introduced in Narayan et al. (2019), accounts for any additional band-dependent systematic scatter that is not accounted for in the measurement error of each observed object. If there are no additional systematics and the measurement errors describe the data perfectly, the model will infer $\sigma_{\text{int}}^{X}$ to be zero. We only consider a CRNL coefficient in the F160W band and set the coefficients for the remaining bands to zero. There is also an additional zeropoint offset $\Delta Z_{X}^{c}$ accounting for changes in zeropoints between Hubble cycles (Narayan et al., 2019; Calamida et al., 2022a; Marinelli et al., 2024) and differences in photometric settings. For at least one Hubble Cycle we must set $\Delta Z_{X}^{c}=0$ to avoid degeneracy with $Z_{X}$. For our analysis we set $\Delta Z_{X}^{25}=0$, then infer $\Delta Z_{X}^{20}$, and $\Delta Z_{X}^{22}$.
In order to infer zeropoints we need to tie our network to a photometric system. In principle this means setting at least one magnitude in a single band to a constant value. For self consistency, we arbitrarily tie the three bright WDs to the Bohlin et al. (2020) CALSPEC system by setting the F475W magnitudes of primary standards, G191B2B, GD153 and GD71 to constant. The primary standards have known calibrated magnitudes $\bar{m}^{X}_{p}$ that are calculated by integrating published SEDs333g191b2b_mod_012.fits, gd153_mod_012.fits and gd71_mod_012.fits from Bohlin et al. (2020) available at: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/(Bohlin et al., 2020) with the HST/WFC3 filter responses. The likelihood is therefore only dependent on the photometry parameters $\bm{\theta_{\text{phot}}^{X}}$ and $\Delta Z_{X}^{c}$ such that
$P\big{(}\hat{m}^{X}_{i,c,p}|\ \bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}=\\ T\big{(}\hat{m}^{X}_{i,c,p}|\ \bar{m}^{X}_{p}-Z_{X}-\Delta Z_{X}^{c}-\alpha_{X}(\bar{m}^{X}_{p}-15),\sqrt{\sigma^{X\hskip 8.5359pt2}_{s,i,c}+\sigma_{\text{int}}^{X\hskip 2.84544pt2}},\nu^{X}\big{)}$ (5)
Figure 4: Example of modelled spectrum for WDFS1206-27. The black lines represent the observed spectra for the source from HST/STIS (left) and SOAR (right). The orange lines represent the inferred SED at the observed wavelengths. The green lines are the inferred cubic splines accounting for residual differences between the model and the data. These splines account for instrumental effects and inadequacies of the model, such as the dust relation or theoretical DA WD template. The blue line represents the modelled SED with the addition of the inferred cubic splines.
3.3 Spectroscopic Modelling
For the spectroscopic modelling, we renormalise the reddened SED $\tilde{F}(\lambda;\bm{\Theta_{\text{WD}}}\big{)}=\tilde{F}(\lambda;\log g,T_{\text{eff}},A_{V},R_{V})$ defined in Equation (1) and convolve it with a Gaussian kernel such that
$f\big{(}\lambda;\bm{\Theta^{s}_{\text{WD}}},\gamma^{j}_{s},\text{FWHM}_{s}^{j}\big{)}=\\ \frac{\tilde{F}(\lambda;\bm{\Theta^{s}_{\text{WD}}})}{4\pi\gamma_{j,s}^{2}}\hskip 2.84544pt\ast\hskip 2.84544ptN\big{(}\lambda|\ 0,\sigma_{R}(\text{FWHM}_{s}^{j})^{2}\big{)}$ (6)
The normalisation factor $\gamma^{j}_{s}$ acts as a nuisance factor to scale the synthetic spectrum to match observed spectrum $j$. The Gaussian kernel was introduced in Narayan et al. (2019) to smooth the synthetic spectrum to be consistent with the observed spectrum’s resolution, where absorption features are less sharp. The standard deviation of the kernel is defined as
$\sigma_{R}(\text{FWHM})=\frac{\text{FWHM}}{R\cdot\sqrt{8\ln 2}}$ (7)
where $R$ is the resolution of the spectrum in Å per spectral unit and FWHM determines the extent of the smoothing.
Further flexibility in the model is needed to correct for wavelength-dependent systematics caused by instrumental or calibration effects. This flexibility can also be thought of as inconsistencies between the model and the data, where either the dust relation or theoretical DA WD template disagrees with what is observed (Czekala et al., 2015). In Narayan et al. (2019) the wavelength-dependent systematics were modelled using a Gaussian Process (GP; Rasmussen & Williams 2005) as an additive function. This approach allowed for maximum flexibility to learn the systematic but was also computationally expensive, involving large multivariate Gaussian covariances with the same dimensionality as the number of spectrum pixels. There was evidence in Narayan et al. (2019) (see Figure 9 of their work) that suggested the inferred function was simple and smooth, allowing us to choose a computationally cheaper method in our hierarchical work.
In our work we choose to model wavelength-dependent spectroscopic systematics using cubic splines. We take inspiration from Mandel et al. (2022) by separating the cubic spline into a matrix $\bm{W}_{s}^{j}(\lambda;\bm{l}_{s}^{j})$ of shape $N_{\text{pix}}\times M$, that is predetermined by the locations of spline knots $\bm{l}_{s}^{j}=\{l^{j,s}_{1},...,l^{j,s}_{M}\}$, and a vector of spline coefficients $\bm{\epsilon}_{s}^{j}$ of shape $M\times 1$, such that
$\epsilon(\lambda;\bm{\epsilon}_{s}^{j})=\bm{W}_{s}^{j}(\lambda;\bm{l}_{s}^{j})\bm{\epsilon}_{s}^{j}$ (8)
We choose to place the $M=10$ spline knots equally between the minimum and maximum wavelength coverage of each observed spectrum $j$. The spline knot locations avoid absorption features and are far enough apart to no add to them. We add the spline to the smoothed synthetic SED to make the modelled synthetic spectrum $S(\lambda;\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})$ using
$S(\lambda;\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})=f\big{(}\lambda;\bm{\Theta^{s}_{\text{WD}}},\gamma^{j}_{s},\text{FWHM}_{s}^{j}\big{)}+\epsilon\big{(}\lambda;\bm{\epsilon}_{s}^{j}\big{)}$ (9)
where we gather all synthetic spectrum nuisance parameters into $\bm{\theta}^{j,s}_{\text{spec}}=(\gamma^{j}_{s},\text{FWHM}_{s}^{j},\bm{\epsilon}_{s}^{j})^{T}$. Figure 4 demonstrates how including the smooth spline corrects systematics in modelled spectra and reduces residuals.
We then construct the likelihood of the WD parameters given the observed spectrum $\bm{\hat{S}^{j}_{s}}$, representing the $j$th observation belonging to WD $s$, as
$P(\bm{\hat{S}}^{j}_{s}|\ \bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}})=\prod_{k=1}^{K_{s}^{j}}N\bigg{(}\hat{S}^{k,j}_{s}\bigg{|}\ S\big{(}\lambda^{s,j}_{k},\bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}}\big{)},\sigma^{j,s\hskip 1.42271pt2}_{k}\bigg{)}$ (10)
We denote the wavelength pixel with $k$, where $\lambda^{s,j}_{0}$ is the first observed wavelength of spectrum $j$ measuring WD $s$ and $\lambda^{s,j}_{K}$ is the last. The observed flux value and corresponding error of the $k$th pixel in the observed spectrum are represented by $\hat{S}^{k,j}_{s}$ and $\sigma^{j,s}_{k}$. We make the assumption that each spectroscopic pixel $\hat{S}^{k,j}_{s}$ is conditionally independent given the WD parameters $\bm{\Theta^{s}_{\text{WD}}}$ and spectrum nuisance parameters $\bm{\theta}^{j,s}_{\text{spec}}$. Importantly the formulation of the spectroscopic modelling is robust to model multiple spectra $j$ per object with different wavelength coverage. In our work this feature is useful for inference using both optical and UV data, but in future work it is trivial to also include near-infrared spectra. For WDs that have multiple optical spectra we choose the spectrum used in Narayan et al. (2019).
3.4 Hierarchical Model
In constructing the hierarchical model, we make a series of assumptions that enable us to treat the random variables as independent. To begin with, we assume that intrinsic WD parameters are independent and that dust parameters are conditionally independent given their respective population hyperparameters. We assume that modelled photometry and modelled spectra for a given WD are conditionally independent given their intrinsic parameters and dust parameters. Each modelled spectrum is assumed to be conditionally independent given its WD parameters and has its own set of independent nuisance parameters. Modelled photometry for a given WD is conditionally independent given its photometric and cycle-dependent systematic parameters. We assume that all bands and cycles have their own set of systematics that are also independent. Finally, all observed photometric and spectroscopic measurement errors are assumed to be independent from each other and model parameters. The relationships are illustrated in the probabilistic graphical model in Figure 3.
We combine the photometric and spectroscopic likelihoods into a combined hierarchical data likelihood for the full network at the same time, such that
$P(\bm{\mathcal{D}}|\ \bm{\Theta})=\\ \hskip 11.38092pt\prod_{s}^{N_{\text{WD}}}\Bigg{[}\prod_{X}^{N_{\text{bands}}}\prod_{c}^{N_{\text{cycles}}}\prod_{i}^{N^{s,X,c}_{\text{Obs}}}P\big{(}\hat{m}^{X}_{i,c,s}\big{|}\ \bm{\Theta^{s}_{\text{WD}}},\mu_{s},\bm{\theta_{\text{phot}}^{X}},\Delta Z_{X}^{c}\big{)}\\ \times\prod_{j}^{N^{s}_{\text{spec}}}P\big{(}\bm{\hat{S}}^{j}_{s}\big{|}\ \bm{\Theta^{s}_{\text{WD}}},\bm{\theta}^{j,s}_{\text{spec}}\big{)}\Bigg{]}$ (11)
where $\bm{\Theta}$ contains all white dwarf, spectroscopic and photometric parameters. We denote all spectroscopic and photometric data by $\bm{\mathcal{D}}$.
We also define an additional likelihood term for the F475W band data $\bm{\mathcal{D}_{p}}$ of the $N_{\text{prim}}=3$ CALSPEC primary standards from Bohlin et al. (2020) as
$P(\bm{\mathcal{D}_{p}}|\ \bm{\Theta})=\\ \prod_{p}^{N_{\text{prim}}}\prod_{c}^{N_{\text{cycles}}}\prod_{i}^{N^{p,\text{F475W},c}_{\text{Obs}}}P\big{(}\hat{m}^{\text{F475W}}_{i,c,p}\big{|}\ \bm{\theta_{\text{phot}}^{\text{F475W}}},\Delta Z_{\text{F475W}}^{c}\big{)}$ (12)
The likelihood terms are combined with the priors on parameters using Bayes’ rule to derive the posterior
$P(\bm{\Theta}|\ \bm{\mathcal{D}})\propto P(\bm{\mathcal{D}}|\ \bm{\Theta})P(\bm{\mathcal{D}_{p}}|\ \bm{\Theta})P(\bm{\Theta})$ (13)
For sampling the hierarchical model we use Hamiltonian Monte Carlo (Duane et al., 1987), a technique that uses gradient information (something that neural network architectures also use) for faster convergence on the posterior. We use the No-U-Turn Sampler (NUTS; Hoffman & Gelman 2011) in numpyro (Bingham et al., 2019; Phan et al., 2019), a probabilistic programming package for automatic differentiation and just-in-time (JIT) compilation built with JAX (Bradbury et al., 2018). Using this software combination on a GPU node, the implementation takes approximately 30 minutes in total to produce four chains of 500 burn-in samples and 500 posterior samples, amounting to an average effective sample size of 1277. This is a significant speed-up compared to the previous implementation of the non-hierarchical model from Narayan et al. (2019) that would take roughly one CPU day per object to converge. When assessing the quality of our chains we diagnose convergence by confirming the Gelman-Rubin $\hat{R}$ statistic is less than 1.02 for each of the 3653 inferred parameters (Gelman & Rubin, 1992; Vehtari et al., 2021). We also find that none of the chains result in divergent transitions (Betancourt et al., 2014; Betancourt, 2017). The implementation of DAmodel can be found at: https://github.com/benboyd97/DAmodel.
3.5 Priors
In our modelling we use non-informative priors $P(\bm{\Theta})$ on our parameters. The priors on intrinsic WD and dust parameters are defined as
$P(\log g^{s})$ $=U(\log g^{s}|\ 7\text{},9.5\text{ })$ (14)
$P(T_{\text{eff}}^{s})$ $=U(T_{\text{eff}}^{s}|\ 15000\text{ K },70000\text{ K })$ (15)
$P(R_{V}^{S}|\ \mu_{RV},\sigma_{RV})$ $=TN(R_{V}^{s}|\ \mu_{RV},\sigma_{RV}^{2},1.2,\infty)$ (16)
$P(A_{V}^{s}|\ \tau)=\begin{cases}\frac{1}{\tau}\exp\Bigg{(}-\frac{A_{V}^{s}}{\tau}\Bigg{)}&\text{if $A^{s}_{V}\geq$ 0}\\ 0&\text{if $A^{s}_{V}<$ 0}\end{cases}$ (17)
[…]
The priors on our population dust parameters when we infer them are
[…]
where $HC(0,1)$ is a Half-Cauchy distribution with a scale of 1. The lower bound on $R_{V}$ of 1.2 is motivated by the Rayleigh scattering limit (Draine, 2003).
The priors on $\mu_{s}$ and photometric parameters are defined as
[…]
$P(\sigma_{\text{int}}^{X})$ $=HC(\sigma_{\text{int}}^{X}|\ 0,1)$ (22)
$P(Z_{X})$ $=N(Z_{X}|\ \mu_{Z}^{X},1^{2})$ (23)
[…]
The $\mu_{Z}^{X}$ parameter is determined by taking the average difference between the measured (Cycle 25 only) and known calibrated magnitudes for the three CALSPEC standards for each band.
The spectroscopic nuisance parameter priors are defined as
[…]
$P(\text{FWHM}^{j}_{s})$ $=TN(\text{FWHM}^{j}_{s}|\ 0,8^{2},0,50)$ (28)
[…]
Table 1: White dwarf and dust parameters inferred using DAmodel. Parameters for stars with an * were inferred from STIS UV spectra as well as ground based optical spectra. Results shown are the averages and standard deviations from our chains. Note that the achromatic offset $\mu$ cannot be treated exactly as distance modulus, as as the absolute magnitude of each WD is not known without imposing a mass-radius-luminosity relation.
Object $T_{\text{eff}}$ $\log g$ $A_{V}$ $R_{V}$ $\mu$
K mag mag
G191B2B $61941\pm 519$ $7.67\pm 0.03$ $0.0\pm 0$ 0.0 $3.1\pm 0.18$ $52.59\pm 0.01$
GD153 $41437\pm 276$ $7.76\pm 0.03$ $0.02\pm 0.0$ $3.18\pm 0.15$ $53.74\pm 0.01$
GD71 $33719\pm 118$ $7.88\pm 0.01$ $0.02\pm 0.0$ $3.24\pm 0.16$ $53.13\pm 0.01$
WDFS0103-00 $64833\pm 3796$ $7.55\pm 0.06$ $0.13\pm 0.01$ $2.94\pm 0.16$ $60.06\pm 0.05$
WDFS0228-08 $22597\pm 483$ $7.96\pm 0.03$ $0.14\pm 0.02$ $3.11\pm 0.17$ $59.17\pm 0.02$
WDFS0248+33 $33646\pm 483$ $7.29\pm 0.06$ $0.34\pm 0.01$ $3.36\pm 0.08$ $58.32\pm 0.02$
WDFS0727+32 $54754\pm 3101$ $7.79\pm 0.05$ $0.16\pm 0.01$ $2.91\pm 0.13$ $58.77\pm 0.05$
WDFS0815+07 $36439\pm 1623$ $7.22\pm 0.06$ $0.09\pm 0.02$ $3.05\pm 0.18$ $60.1\pm 0.05$
WDFS1024-00 $36012\pm 1350$ $7.79\pm 0.13$ $0.23\pm 0.02$ $3.06\pm 0.15$ $59.09\pm 0.05$
WDFS1110-17 $52672\pm 1921$ $7.8\pm 0.04$ $0.19\pm 0.01$ $3.25\pm 0.11$ $58.56\pm 0.03$
WDFS1111+39 $56759\pm 3481$ $7.84\pm 0.07$ $0.02\pm 0.01$ $3.1\pm 0.18$ $59.39\pm 0.05$
WDFS1206+02 $24243\pm 272$ $7.96\pm 0.02$ $0.08\pm 0.01$ $3.12\pm 0.16$ $58.22\pm 0.01$
WDFS1214+45 $36003\pm 438$ $7.97\pm 0.03$ $0.05\pm 0.01$ $3.11\pm 0.16$ $58.18\pm 0.01$
WDFS1302+10 $46154\pm 887$ $7.87\pm 0.02$ $0.1\pm 0.01$ $2.87\pm 0.11$ $57.71\pm 0.02$
WDFS1314-03 $43685\pm 2696$ $7.79\pm 0.07$ $0.11\pm 0.02$ $3.24\pm 0.16$ $59.69\pm 0.05$
WDFS1514+00 $29853\pm 74$ $7.88\pm 0.01$ $0.18\pm 0.0$ $3.45\pm 0.08$ $55.6\pm 0.01$
WDFS1557+55 $66877\pm 1942$ $7.57\pm 0.05$ $0.05\pm 0.01$ $2.95\pm 0.17$ $58.57\pm 0.02$
WDFS1638+00 $65645\pm 3408$ $7.63\pm 0.13$ $0.23\pm 0.01$ $3.18\pm 0.14$ $59.7\pm 0.04$
WDFS1814+78 $32019\pm 144$ $7.91\pm 0.02$ $0.05\pm 0.0$ $3.06\pm 0.14$ $56.75\pm 0.01$
WDFS2101-05 $29696\pm 358$ $7.86\pm 0.05$ $0.17\pm 0.01$ $3.22\pm 0.13$ $58.54\pm 0.02$
WDFS2329+00 $21729\pm 332$ $7.99\pm 0.03$ $0.19\pm 0.02$ $3.13\pm 0.12$ $57.36\pm 0.02$
WDFS2351+37 $47766\pm 1354$ $7.93\pm 0.06$ $0.37\pm 0.01$ $3.12\pm 0.07$ $58.45\pm 0.03$
WDFS0122-30 $36474\pm 628$ $7.66\pm 0.05$ $0.08\pm 0.01$ $3.08\pm 0.16$ $58.86\pm 0.02$
WDFS0238-36 $24094\pm 225$ $7.88\pm 0.04$ $0.21\pm 0.01$ $3.13\pm 0.1$ $57.48\pm 0.01$
WDFS0458-56 $31128\pm 218$ $7.79\pm 0.04$ $0.04\pm 0.01$ $3.1\pm 0.17$ $57.9\pm 0.01$
WDFS0541-19 $21144\pm 195$ $7.88\pm 0.03$ $0.1\pm 0.01$ $3.25\pm 0.16$ $57.54\pm 0.01$
WDFS0639-57 $54352\pm 1811$ $7.71\pm 0.07$ $0.19\pm 0.01$ $3.61\pm 0.12$ $58.92\pm 0.03$
WDFS0956-38 $19791\pm 79$ $7.87\pm 0.01$ $0.12\pm 0.01$ $3.38\pm 0.14$ $56.97\pm 0.01$
WDFS1055-36 $30682\pm 199$ $7.93\pm 0.03$ $0.15\pm 0.01$ $3.31\pm 0.11$ $58.0\pm 0.01$
WDFS1434-28 $20790\pm 128$ $7.87\pm 0.03$ $0.22\pm 0.01$ $3.33\pm 0.11$ $57.07\pm 0.01$
WDFS1837-70 $19975\pm 91$ $7.89\pm 0.02$ $0.14\pm 0.01$ $3.02\pm 0.12$ $56.88\pm 0.01$
WDFS1930-52 $38498\pm 632$ $7.79\pm 0.04$ $0.17\pm 0.01$ $3.33\pm 0.11$ $57.86\pm 0.02$
WDFS1206-27 $35884\pm 361$ $7.8\pm 0.04$ $0.15\pm 0.01$ $3.35\pm 0.11$ $56.77\pm 0.01$
WDFS1535-77 $62316\pm 1428$ $8.97\pm 0.03$ $0.07\pm 0.0$ $2.85\pm 0.15$ $57.54\pm 0.02$
WDFS2317-29 $25063\pm 224$ $7.82\pm 0.03$ $0.08\pm 0.01$ $2.83\pm 0.16$ $57.97\pm 0.01$
Figure 5: Ratio of SEDs inferred with and without using STIS UV spectra. Both SEDs were also inferred using optical spectra and HST/WFC3 photometry. The red line shows the average flux ratio as a function of wavelength, while the grey shows the range of ratios across the network of 35 stars. The sharp peak around 1200 Å corresponds to the Lyman-$\alpha$ absorption feature that is on average 12% stronger after the introduction of the STIS UV spectra.
Figure 6: Comparing inferred parameters with and without STIS UV spectra. Inferences in top row use a constant $R_{V}=3.1$ for all objects, whilst the bottom row experiments jointly inferred the $R_{V}$ dust ratio for each object.
4 Results and Discussion
In this section we present our calibration results using our hierarchical Bayesian model. The main results presented are those that use all the optical ground-based spectra, the STIS UV space-based spectra for the 19 WDs that have coverage, and the six HST/WFC3 bands ranging the UV to the NIR. The published results are also those that infer $R^{s}_{V}$ for the WDs, rather than keeping it constant as in previous analyses. The prior used in the $R^{s}_{V}$ inference is a truncated normal prior $R^{s}_{V}\sim TN(3.1,0.18,1.2,\infty)$ motivated by Schlafly et al. (2016), where $\sim$ denotes “is sampled from”. In these main results we do not infer dust population parameters to avoid the effects of shrinkage, where inferences of object-level parameters are skewed towards the inferred population average. Our results are on the CALSPEC Bohlin et al. (2020) magnitude system, tied at the F475W band, but may easily be converted to the Bohlin (2014) system using the zeropoints presented in Table 4. The SEDs for all 35 WDs in the network are defined across the entire range of the Gordon et al. (2023) dust relation, spanning from 912 Å to 32 $\mu$m, and are publicly accessible at: https://zenodo.org/records/14339961
4.1 Inferred Parameters and SEDs
The results of our inferred intrinsic and dust parameters for the 35 WDs are presented in Table 1. These five parameters per object are all that is necessary to forward model the calibrated SEDs. We illustrate the derived SEDs for the 32 faint standards, as well as the three CALSPEC primary standards, in Figure 12.
[…]
We can also compare the new SEDs with those published in Axelrod et al. (2023) in Figure 13. The differences in SEDs are due to numerous differences in methodology and modelling. The smooth decrease in average flux we see in the UV, including the bump at 2175 Å, are due to the differences in dust relation when we transition from Fitzpatrick (1999) to Gordon et al. (2023) dust relations. In our work we jointly infer $R_{V}$ for each WD, whereas in previous works this was kept to a constant $R_{V}=3.1$. The sharp Balmer series flux increase of up to 7% in the optical is likely due to the introduction of the STIS UV data influencing inferred $A_{V}$ and $T_{\text{eff}}$. In Figure 14 we ignore the STIS data and choose the same modelling decisions as Axelrod et al. (2023) finding we infer parameters that are largely consistent with the previous analysis. In Section 4.5 we further explore the influence of using the Gordon et al. (2023) dust relation over the Fitzpatrick (1999) dust relation. Even after these modelling choices we would still expect to see different inferred SEDs between our work and the previous analysis, due to the changes in the updated theoretical DA WD template, the hierarchical nature of our model and the spectroscopic-informed priors in Axelrod et al. (2023).
In our main results we infer $R^{s}_{V}$ for each DA WD using a truncated normal prior with $\mu_{RV}=3.1$ and $\sigma_{RV}=0.18$, physically motivated by Schlafly et al. (2016). The distribution of inferred $R^{s}_{V}$ parameters do not deviate significantly from this prior using both Gordon et al. (2023) and Fitzpatrick (1999) dust relations. We find the average extinction $\tau$ does not deviate significantly above the prior value of 0.1 mag. We find WDFS2351+37 has a strong extinction of $A_{V}^{s}=0.372\pm 0.009$ mag ($R_{V}^{s}=3.12\pm 0.07$), which is indicative of anomalous reddening of this source. This finding is consistent with Axelrod et al. (2023) and independent analysis by Bohlin et al. (2024). We also find strong extinction of $A^{s}_{V}=0.34\pm 0.01$ mag ($R_{V}=3.36\pm 0.08$) for WDFS0248+33, which is again consistent with Axelrod et al. (2023).
[…]
4.2 Spectroscopic Spline Residuals
An important part of our methodology involves inferring a cubic spline to account for residual differences between the observed spectroscopic data and the modelled spectral energy distribution. In Figure 4 we show the model fit to WDFS1206-27 as an example. These splines may represent wavelength-dependent systematics in the ground or space-based spectrograph. Alternatively, the splines may be suggesting inadequacies in the modelling, where the theoretical SED model or the dust relation are unable to express the data. In Narayan et al. (2019) these residuals were modelled with a GP, but the functions themselves were found to be simple and smooth. In this work we instead chose to use a cubic spline with 10 equally spaced knots. We experimented with multiplying the splines in flux-space, but found this gave unreliable results, allowing absorption features to be artificially strengthened at different wavelengths, leading to differences in inferred parameters. Instead we settled on the adding the spline in flux-space, as this is still physically motivated by instrumentals effects and avoids degeneracies. In Figure 20 we plot the relative strength of each inferred spline for the STIS UV and optical spectra.
The ground-based spectroscopic data may have been influenced by systematics introduced when calibrating to a set of standards stars. These standards themselves are only accurate to a few percent, whereas we require sub-percent accuracy. For this reason it was important to include extra flexibility to allow the model to deviate from the data. Our optical splines in the right panel of Figure 20 show larger deviations from the model at the bluer (<4200 Å) edges of the spectra. These differences were in agreement with what was found in Narayan et al. (2019) using GPs and are thought to be due to biased flux corrections at the edges of the spectrograph, where the throughput drops sharply.
Although STIS UV space-based spectra are less impacted by calibration systematics compared to optical spectra, it remains crucial to address other instrumental effects. In Figure 20, we see the strength of the inferred STIS UV splines decrease as a function of wavelength. The residuals are largely negative suggesting there may be an underestimation of flux in the STIS UV spectra. The STIS MAMA detectors dark current performance has degraded steadily in-flight and the dark current subtraction by the “On the Fly Reprocessing (OTFR)” pipeline may underestimate the dark level, leading to an additional background. As STIS uses two separate channels (FUV: 1150-1700 Å and NUV: 1570-3180 Å) and the final spectrum is reconstructed from both, the background level can exhibit a wavelength dependence as we see in the calibration splines, which increase in amplitude below 1700 Å. Additionally, our faint DA standards require longer exposures than the bright standards used to calibrate STIS, and estimate the dark current rate, and a small misestimation of the dark rate in the FUV MAMA would be sufficient to cause an additive flux excess in the final spectrum. However, we cannot uniquely identify the STIS dark current as the cause for the wavelength dependence seen in the STIS spectrum calibration splines, and in particular, sunlight reflected off the Earth or Moon can scatter onto the STIS detectors, and is dependent on the angle between each standard and the bright limb of the Earth/Moon, while scattered zodiacal light depend on helio-ecliptic coordinates [^4] Information on STIS instrumental systematics: https://www.stsci.edu/hst/instrumentation/stis/performance/background-noise. The intensity of scattered light is expected to drop as a function of wavelength, which is opposite to the trend we see in the STIS data. However, even if the intensity of scattered light drops in the UV, if the correction for scattered light is underestimated in the UV, it could yield a chromatic effect similar to what is seen in our splines.
Figure 7: Final residual results achieved when using our model. Results show unweighted mean and RMS residuals for the 32 faint DA WDs, excluding the three bright CALSPEC standards. Scatter points are arranged in order of brightness, ending with the dimmest object. The results were found using the STIS UV spectra, optical spectra and the HST/WFC3 photometry. We use the Gordon et al. (2023) dust relation and infer $R_{V}$ jointly per object.
The deviations between model and STIS UV spectra exhibited by the splines may also be hinting that the model needs to be improved at short wavelengths. The theoretical intrinsic DA WDs templates have recently been updated with improved UV modelling, but it is possible that more work needs to be done. Alternatively, there is a good chance that the differences may be due to the chosen dust relation. Dust relations like Gordon et al. (2023) and Fitzpatrick (1999) are averages of how extinction is observed to change as a function of $A_{V}$ and $R_{V}$. We find that the two strongest STIS splines belong to the object with the largest inferred $A^{s}_{V}=0.37$ and the object with the largest inferred $R^{s}_{V}=3.61$. If the dust relations are averages, it may be possible that these extreme regions of the $R_{V}-A_{V}$ space are poorly sampled leading to extinction curves that are at odds with our observed data. Besides these extremes, we find no obvious correlation between the dust parameters and inferred splines.
We are confident that the splines are not influencing the reliability of our published SEDs. Instead the splines are allowing us to utilise our data in a flexible manner that does not place too much emphasis on that relative calibration of each wavelength pixel. If the splines were influencing the inferred intrinsic parameters, we would see them trace the shapes of the Balmer series, either adding or removing signal. If the STIS splines were adding to extinction, we would expect them to trace the 2175 Å bump. Conversely, we find our inferred splines to be smooth and to not exhibit traits that could be misinterpreted as additional signal coming from our model parameters. This conclusion is further supported by the observation that the inferred parameters in Figure 6 do not exhibit the extreme change one would expect if they were artificially influenced by the STIS spline.
Figure 8: Comparison between photometric residuals of this work and Axelrod et al. (2023). The blue and green experiments utilised the STIS UV spectra for the 19 objects that had it, whilst the others experiments only used optical spectra. The blue and orange experiments used a truncated normal prior distribution to infer $R_{V}$ whilst the other experiments assumed it to be a constant 3.1. RMS residuals are unweighted including only the 32 faint DA WD stars and not the three CALSPEC standards.
Table 2: Inferred HST/WFC3 photometric systematics obtained in our final results.
Parameter F275W F336W F475W F625W F775W F160W
$Z_{X}$ (mag) 23.978$\pm$0.002 24.544$\pm$0.001 25.572$\pm$0.001 25.407$\pm$0.001 24.732$\pm$0.001 25.789$\pm$0.002
$\Delta Z_{X}^{20}$ (mag) - -0.004$\pm$0.003 -0.02$\pm$0.003 -0.025$\pm$0.003 -0.02$\pm$0.004 -0.007$\pm$0.006
$\Delta Z_{X}^{22}$ (mag) 0.072$\pm$0.001 0.031$\pm$0.001 -0.007$\pm$0.001 -0.014$\pm$0.001 -0.028$\pm$0.001 0.002$\pm$0.001
$\sigma_{\text{int}}^{X}$ (mmag) 2.57$\pm$0.56 2.25$\pm$0.49 2.78$\pm$0.34 2.13$\pm$0.56 0.5$\pm$0.36 6.79$\pm$0.65
[…]
$\alpha_{X}$ (mmag/mag) - - - - - -3.19$\pm$0.31
Notes: $Z_{X}$ represents the zeropoint magnitude at Cycle 25 (average MJD=57588.79 days) in a given band $X$. Zeropoints are calculated from our aperture radii (7.5 pixels for UVIS and 5 pixels for the IR) to infinity and are in AB mags. In Appendix F we compare inferred zeropoints to CALSPEC and relative shifts derived in Axelrod et al. (2023). $\Delta Z_{X}^{20}$ (average MJD=56452.94 days) and $\Delta Z_{X}^{22}$ (average MJD=57076.11 days) the change in zeropoint in each band in Cycle 20 and Cycle 22. We have no constraint on $\Delta Z_{\text{F275W}}^{20}$ as we only began F275W observations from Cycle 22. $\sigma_{\text{int}}^{X}$ and $\nu^{X}$ represent the student-T scale and dispersion parameters for each band. $\alpha_{\text{X}}$ represents CRNL parameter inferred only in the NIR F160W band. Estimates and uncertainties are the means and standard deviations of our chains, with the exception of $\nu^{\text{F160W}}$, that has a non-Gaussian posterior, so is presented as a median with the 16th and 84th percentiles.
Table 3: HST/WFC3 synthetic photometry and residuals from our final results. Units are in AB magnitudes and photometry for stars with an * was inferred using STIS UV spectra as well as ground-based optical spectra.
Object F275W F336W F475W F625W F775W F160W
Synth. Resid. Synth. Resid. Synth. Resid. Synth. Resid. Synth. Resid. Synth. Resid.
G191B2B 10.481 0.0 10.879 -0.002 11.488 -0.001 12.018 0.0 12.433 0.002 13.866 0.001
GD153 12.194 -0.0 12.554 0.0 13.086 0.0 13.586 0.0 13.987 -0.0 15.395 -0.001
GD71 11.979 -0.0 12.319 0.0 12.785 -0.001 13.264 0.002 13.657 -0.0 15.048 -0.002
WDFS0103-00 18.18 0.006 18.517 -0.007 19.07 -0.002 19.556 0.001 19.942 0.008 21.318 -0.002
WDFS0228-08 19.509 -0.0 19.698 0.003 19.805 -0.002 20.158 -0.001 20.483 0.003 21.708 -0.014
WDFS0248+33 17.817 0.002 18.032 -0.008 18.354 0.003 18.736 -0.003 19.061 0.001 20.307 -0.005
WDFS0727+32 17.151 0.003 17.462 -0.006 17.979 0.001 18.445 -0.001 18.821 0.001 20.182 -0.002
WDFS0815+07 18.937 0.004 19.252 -0.003 19.705 -0.002 20.175 -0.003 20.557 0.008 21.923 -0.004
WDFS1024-00 18.248 0.003 18.497 -0.0 18.89 -0.002 19.303 0.003 19.651 -0.001 20.955 -0.004
WDFS1110-17 17.032 -0.001 17.34 -0.001 17.847 0.007 18.304 -0.003 18.674 0.0 20.017 0.003
WDFS1111+39 17.435 -0.001 17.817 -0.002 18.406 0.002 18.923 0.003 19.332 -0.002 20.753 0.004
WDFS1206+02 18.232 -0.002 18.468 0.003 18.661 -0.003 19.049 0.0 19.395 0.001 20.667 -0.001
WDFS1214+45 16.931 -0.0 17.266 0.001 17.749 -0.001 18.224 -0.001 18.614 0.0 20.001 -0.0
WDFS1302+10 16.178 0.0 16.507 0.0 17.025 -0.001 17.502 -0.001 17.888 0.001 19.269 0.001
WDFS1314-03 18.251 -0.003 18.578 0.002 19.084 0.003 19.557 -0.0 19.94 0.0 21.311 -0.022
WDFS1514+00 15.104 -0.003 15.369 0.006 15.697 -0.001 16.107 0.001 16.458 -0.001 17.754 0.003
WDFS1557+55 16.487 0.003 16.865 -0.003 17.459 -0.002 17.974 0.005 18.379 -0.006 19.791 0.006
WDFS1638+00 18.003 0.003 18.305 -0.004 18.82 0.005 19.273 -0.003 19.638 0.007 20.969 -0.011
WDFS1814+78 15.782 -0.0 16.105 0.002 16.532 -0.001 16.994 -0.001 17.378 -0.0 18.751 0.002
WDFS2101-05 18.06 -0.001 18.32 -0.001 18.641 0.002 19.05 0.0 19.402 -0.003 20.702 -0.002
WDFS2329+00 17.937 -0.003 18.088 0.004 18.136 0.009 18.461 -0.002 18.768 -0.008 19.955 0.005
WDFS2351+37 17.434 0.006 17.652 -0.005 18.059 0.003 18.447 -0.0 18.773 -0.001 20.035 0.002
WDFS0122-30 17.661 0.001 17.981 -0.0 18.447 0.001 18.916 -0.005 19.299 0.005 20.671 -0.004
WDFS0238-36 17.782 -0.001 17.959 -0.0 18.082 0.001 18.427 0.0 18.743 -0.002 19.954 0.002
WDFS0458-56 17.015 -0.001 17.335 0.002 17.743 -0.001 18.204 0.001 18.588 -0.003 19.956 0.008
WDFS0541-19 18.014 -0.003 18.199 0.003 18.262 0.002 18.615 -0.002 18.942 0.002 20.16 -0.003
WDFS0639-57 17.313 0.0 17.639 0.0 18.162 0.004 18.628 -0.001 19.002 -0.002 20.349 -0.006
WDFS0956-38 17.691 -0.002 17.844 0.002 17.837 0.013 18.168 -0.0 18.481 -0.0 19.666 -0.01
WDFS1055-36 17.362 -0.001 17.64 0.0 17.998 0.003 18.42 -0.004 18.779 -0.002 20.096 0.002
WDFS1434-28 17.829 -0.0 17.962 0.001 17.96 -0.003 18.271 0.003 18.569 -0.001 19.729 -0.004
WDFS1837-70 17.635 -0.001 17.774 0.004 17.76 -0.001 18.083 -0.003 18.392 0.003 19.573 -0.002
WDFS1930-52 16.722 -0.001 17.018 0.002 17.472 0.001 17.917 -0.001 18.284 0.0 19.623 -0.002
WDFS1206-27 15.729 -0.001 16.027 0.001 16.465 -0.0 16.91 0.002 17.278 -0.001 18.619 -0.002
WDFS1535-77 15.588 0.002 15.954 0.002 16.542 0.0 17.042 -0.003 17.441 -0.0 18.854 0.003
WDFS2317-29 17.886 0.002 18.129 -0.001 18.34 -0.002 18.739 -0.002 19.089 0.0 20.373 0.012
4.3 HST/WFC3 Photometric Residuals and Systematics
In our work we achieve the lowest residuals yet when using our modelled SEDs to predict the observed data, illustrated in Figure 7. We use the 2000 samples of WD and dust parameters to forward model 2000 randomly sampled SEDs per object. To obtain the synthetic photometry we integrate the SED samples using Equation (3) over the six HST/WFC3 bandpasses described in Calamida et al. (2019, 2022b). We then define our synthetic photometry and uncertainty as the mean and standard deviation of the 2000 sampled magnitudes in each band for each object. We use the average of each inferred systematic (presented in Table 2) to correct for instrumental effects to derive the calibrated observed photometry. A detailed breakdown of the residual and synthetic photometry of each object is shown in Table 3.
For the set of 32 faint standards, the average RMS from the UV to the NIR was 3.9 mmag, with an average bias of -0.06 mmag, using our synthetic SEDs and inferred systematics. Excluding the NIR band, the average RMS drops to 3.1 mmag, closer to the quantum efficiency (QE) limit. In Figure 8, we compare these residuals (shown in blue) with those obtained by Axelrod et al. (2023) (indicated by black stars), highlighting improvement across all bands. It is clear from the plot that these improvements are largely as a result of allowing $R^{s}_{V}$ for each band to be inferred, when in previous analyses this was kept constant. We find that switching from a constant $R_{V}=3.1$ to an free $R_{V}$ according to a truncated normal prior improves results significantly in the NIR F160W and UV F275W. This is predictable as $R_{V}$ is unlikely to be exactly 3.1 for the whole sample and differences in $R_{V}$ are more apparent at these wavelengths. From the plot it is also evident that adding STIS UV spectra data does not improve the RMS photometric residuals, however, we still consider these our main results as they are using the most information to determine the SEDs.
We find small amounts of shrinkage when swapping from the normal $R_{V}$ prior with set mean and variance to the population prior where these are jointly inferred. Due to the additional parameters we find that the population $\mu_{RV}$ reduces the overall residual RMS to 3.6 mmag. We exclude these versions of the SEDs from our final results, as the selection effect heavily influences these population parameters, and we aim to avoid over-fitting the data. Results of these tests and other experiments are presented in Table 5.
The procedure in Narayan et al. (2019) described a two-step process where first photometric measurements and offsets are inferred using a HBM, then these measurements are combined with spectra to serially infer WD parameters. As an investigation into our hierarchical results we serially infer WD parameters using a similar procedure to that described in Narayan et al. (2019) and find again there were no significant differences in inferred WD SED parameters. The differences in the results of the two approaches appeared in the inferred photometric systematics of the HST/WFC3 data. This difference is to be expected as our hierarchical method effectively changes the prior on $m^{X}_{s}$, that would have otherwise been flat in the two-step process, by introducing spectroscopic information. This effect was part of our motivation for using a fully hierarchical model, to allow for more precise inference of systematics by modelling all the data at once.
Figure 9: Photometric residuals when comparing our synthetic DES photometry with real DES observations (Abbott et al., 2021) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.
In Table 2 we present our hierarchically inferred HST/WFC3 photometric systematics for each of the six bands. Zeropoints are calculated from our aperture radii (7.5 pixels for UVIS and 5 pixels for the IR) to infinity. The inferred $Z_{X}$ parameters correspond to UVIS1 and IR zeropoints at Cycle 25 (average MJD=57588.79 days) and are in 1% agreement with those calculated in Calamida et al. (2022a) for the same aperture radius and epoch. In Appendix F, we compare zeropoint shifts in our work and the work of Axelrod et al. (2023) with respect to the CALSPEC Bohlin (2014) and Bohlin et al. (2020) systems. We find good agreement with Axelrod et al. (2023), where we infer decreasing zeropoint shifts as function of wavelength ranging from 11 mmag in the UV to -15 mmag in the NIR for Bohlin et al. (2020).
The inferred $\Delta Z_{X}^{20}$ and $\Delta Z_{X}^{22}$ are the deviations away from the Cycle 25 zeropoint in Cycle 20 (average MJD = 56452.94 days) and Cycle 22 (average MJD = 57076.11 days) respectively. We expect the $\Delta Z_{X}^{c}$ values to be negative if only due to a decrease of the WFC3 UVIS or IR detector sensitivity, as shown by Calamida et al. (2022a) and Marinelli et al. (2024). The percentage in sensitivity decrease of each band is parametrised by a linear slope as a function of time. Since we gather our observations into discrete cycles, and the observations span a limited amount of time (< 5 years per star), it is challenging to compare with the official continuous slope. Despite this, we take averages of the percentage decrease evaluated at our observation times, finding that the $\Delta Z_{X}^{20}$ changes between Cycle 20 and Cycle 25 were within error of those calculated in Calamida et al. (2022a) and Marinelli et al. (2024). Our work infers a $\Delta Z_{\text{F336W}}^{20}$ consistent with zero which is also consistent with the literature.
Cycle 20 observations were collected by using UVIS1 sub-arrays and are on the UVIS1 photometric system (Narayan et al., 2016), while Cycle 22 and Cycle 25 observations were instead collected by using UVIS2 sub-arrays. Cycle 22 images were then re-processed by omitting the default flux conversion performed by HST/WFC3 calibration pipeline (calwf3), resulting in this photometry being on the UVIS2 photometric system (Calamida et al., 2019). Cycle 25 observations were not re-processed and the resulting photometry is on the UVIS1 photometric system (Axelrod et al., 2023). These differences in photometric system do not influence our analysis, since the three primary DA WDs were observed in all cycles at multiple epochs allowing us to determine our own zeropoints and systematics.
Our change in zeropoints for Cycle 22 are dominated by the effect of converting between the UVIS1 and UVIS2 photometric systems, rather than by changes in sensitivity, hence why some changes are positive. For instance, the difference between Cycle 25 and Cycle 22 zeropoints for the F275W is $\Delta Z_{\text{F275W}}^{22}\approx$ +0.07 mag, comparable to the QE difference between the UVIS1 and UVIS2 detectors for this filter, with the UVIS2 chip being more sensitive at bluer wavelengths [^5] HST/WFC3 CCD characteristics and performance:
https://tinyurl.com/wfc3QE
. Note that this does not apply to the F160W filter, since there is only one WFC3 IR detector, so $\Delta Z_{\text{F160W}}^{20}$ and $\Delta Z_{\text{F160W}}^{22}$ are in this case mostly due to the change in sensitivity of WFC3-IR. We can compare the inferred changes zeropoints between Cycle 20 and Cycle 22 to those from Narayan et al. (2019) that used the same dataset. Comparisons find consistency in the difference between the $\Delta Z_{X}^{20}$ and $\Delta Z_{X}^{22}$ in our work and the shifts measured in Narayan et al. (2019).
An important systematic to consider is the count-rate nonlinearity in the F160W band. In our analysis, we infer the CRNL to be $\alpha_{\text{F160W}}$ = -3.19 $\pm$ 0.31 mmag/mag. This constraint is within $1\sigma$ of both the published CRNL value of -3.12 $\pm$ 0.32 mmag/mag from Bohlin & Deustua (2019) and the combined result of -3.0 $\pm$ 0.24 mmag/mag from Riess et al. (2019). The result is also close to the post hoc result of -2.6$\pm$0.52 mmag/mag from Narayan et al. (2019). The constraint, however, is in disagreement with the result of -1.74 $\pm$ 0.32 mmag/mag from Axelrod et al. (2023), that used the same WD sample as this work. Interestingly, if we change our modelling decisions to match those of Axelrod et al. (2023), by using Fitzpatrick (1999), a constant $R_{V}=3.1$ and ignoring the STIS UV data, we obtain $\alpha_{\text{F160W}}$ = -2.12 $\pm$ 0.30 mmag/mag. This result is a strong suggestion that the inferred value of the NIR CRNL is sensitive to dust modelling assumptions. In inference, our model is determining which effects are from natural sources such as dust and which are from instrumental systematics. If we change the definition of what is physically possible, such as fixing or changing dust relations, it will then impact the extent to which effects are inferred to be coming from systematics. The inferred CRNL’s dependency on dust is also supported by correlations shown in Figure 17.
The inferred student-T dispersion $\sigma^{X}_{\text{int}}$ and degrees of freedom $\nu^{X}$ parameters were in close agreement with those inferred by Narayan et al. (2019) from the UV to the optical. The F160W parameter degree of freedom parameter $\nu^{\text{F160W}}$ was higher in our work, suggesting a normal distribution may have been sufficient, likely due to the better handling of the CRNL. The size of $\sigma^{X}_{\text{int}}$ inferred is comparable to our residual RMS in each band. This suggests that we are close to the limit of what can be achieved with this data and would need to address the instrumental systematics causing the inferred additional dispersion to achieve more accurate results.
4.4 Photometric Residuals with other surveys
We demonstrate the use of our resulting SEDs to produce synthetic photometry for five common surveys. To obtain the synthetic photometry for each survey we repeat the same procedure as we describe at the beginning of Section 4.3, but instead integrate over the bandpasses for each survey. We obtain the filter bandpasses from the Spanish Virtual Observatory (SVO) Filter Profile Service (Rodrigo et al., 2012; Rodrigo & Solano, 2020). For surveys like Gaia, we must then do an extra step of converting our synthetic photometry from AB to Vega magnitude units. Only subsets of the 32 faint stars are observed by each survey, but it is still useful to look at the residual trends. The observations for each of the surveys are completely independent of the training of our model so we naturally expect to see larger scatter and biases than we would in the HST photometry. For detail on the observed photometric data obtained for each survey refer to Narayan et al. (2019) and Axelrod et al. (2023), where the same comparisons are made using their SEDs.
Figure 9 shows the residual scatter for seven of our DA WDs in DES DR2 that were observed with DECam (Abbott et al., 2021). We see residual bias in $g$, $r$ and $i$ bands between 12 and 13 mmag, which then reduces to 4 mmag in the $z$ band. We show corresponding results in Figure 21 using SDSS (Holberg & Bergeron, 2006), Figure 22 using PS1 DR2 (Flewelling et al., 2020), Figure 23 using Gaia DR3 (Gaia Collaboration et al., 2023) and Figure 24 using DECaLS (Schlegel et al., 2021) data. Residual bias with other surveys is highly dependent on the magnitude systems used in calibration of both our work and the surveys themselves. In our method we set the primary standard magnitudes as constants in one band according to Bohlin et al. (2020) CALSPEC. We find that when we set the magnitudes to constant in all six bands according to Bohlin (2014) CALSPEC, residual scatter with other surveys decreases in the optical but increases towards the NIR.
These comparisons between our synthetic photometry and observed photometry from other surveys are necessary for cross-calibration when combining datasets. Our all-sky faint standards will allow various surveys to use a single self-consistent flux scale. In Section 5 we discuss efforts that are already using our inferred SEDs in the cross-calibration of datasets.
Figure 10: The ratio of Gordon et al. (2023) dust relation with Fitzpatrick (1999) dust relation, evaluated at different $A_{V}$ and $R_{V}$ values. We see moderate discrepancies in the UV and strong discrepancies in the NIR.
Figure 11: Comparing inferred parameters when using Gordon et al. (2023) dust relation and Fitzpatrick (1999) dust relation.
4.5 Using G23 Over F99 Dust Relation
An important update to our modelling of the DA WDs was the adoption of the Gordon et al. (2023) dust relation over the Fitzpatrick (1999) relation. As illustrated in Figure 10, the main differences between the relations appear in the UV and NIR wavelengths, which are critical regimes for robust inference of $A_{V}$ and $R_{V}$. It is important that our models extrapolate well to wavelengths where we do not have data coverage, which is heavily reliant on the accuracy of the dust relation.
In Figure 11 we compare the inferred parameters when using both the Fitzpatrick (1999) and Gordon et al. (2023) dust relations with a truncated normal $R_{V}$ prior whilst utilising the STIS UV spectra. There is good agreement for $\log g^{s}$ surface gravity inference when using the different dust relations. Switching to Gordon et al. (2023) results in a 7% decrease in $A_{V}$, causing the warmer DA WDs to have lower inferred effective temperatures. The two WDs that deviate the most had STIS UV coverage, suggesting there was less extinction at these wavelengths, decreasing the inferred intrinsic strength of the Lyman-$\alpha$ absorption features. Switching to the updated relation results in $R_{V}^{s}$ inferences that differ on average 2% greater, hinting a slight change in definition, although the uncertainties on this parameter are still large due to the lack of spectroscopic coverage in the NIR. The $R^{s}_{V}$ of WDFS0639-57 increases significantly from 3.22 $\pm$ 0.13 to 3.61 $\pm$ 0.10 when switching form Fitzpatrick (1999) to Gordon et al. (2023), again emphasising a change in definition of the dust relations at parameter extremes. During our population analysis we see large differences in the inferred $\mu_{R_{V}}$, increasing from 3.01 $\pm$ 0.10 to 3.42 $\pm$ 0.10 when switching to the newer relation, however, as previously stated, these results are still influenced heavily by selection effects and small sample sizes.
In Table 5 we can study the effect switching dust relation has on HST/WFC3 residuals. We see using Gordon et al. (2023) dust relation results in a lower average residual bias and RMS scatter across the six bands. The largest improvement was in the F275W UV band where both bias and scatter dropped by roughly 1 mmag. Interestingly, results were marginally worse in the F475W and F625W bands when using Gordon et al. (2023), although these results are already close to QE.
Dust relations are averages so will never completely agree with our observed data. In Section 5 we discuss the possibility of extending our model to simultaneously infer its own dust relation.
5 Future Work
It is a priority to obtain NIR spectra for our faint DA WD standards in the near future by using an instrument such as NIRSpec on JWST (Böker et al., 2022). Note that a sub-sample of these stars are already included in the official JWST calibration program (see proposal 6605, PI Gordon). Adding spectroscopic coverage at IR wavelenghts would allow for more robust inferences of $R_{V}$ and further reduce photometric residuals in the F160W band. Modelling the NIR spectra would be a trivial extension to the model, as it was for the STIS UV spectra, where we simply evaluate the model at a different set of wavelengths and include additional cubic splines. We are also eager to obtain further STIS UV coverage for the remaining 13 DA WDs that do not have it, indicated in Table 1 and Table 3 by the WDs without an asterisk. Obtaining this panchromatic spectroscopic coverage will improve the network’s ability to be NIR standards for JWST itself, as well as Roman and the LSST $y/z$ bands.
The scalable implementation of the model allows for the possibility of expanding the network to more objects. Natural targets to join the network would be the 11 DA WDs observed by Rubin et al. (2022a) using HST/STIS. The 17 cooler DA WDs added to CALSPEC by Elms et al. (2024) could also be a extension to the network providing we update our theoretical intrinsic DA WD templates at these temperatures. Hierarchically modelling this extended sample of DA WDs would further improve internal HST/WFC3 systematic mitigation, whilst having higher statistics would allow future surveys to improve their flux calibration.
Aside from the purpose of calibration, our novel statistical framework opens doors to population analysis on larger samples of DA WDs. There are nearly 2000 DESI DA WDs (Manser et al., 2024) and a further 500 DA WDs in Gaia (O’Brien et al., 2024), with both sets having spectroscopic and photometric coverage. A volume limited subset of these samples could be analysed by our framework to provide more meaningful statistics on population dust parameters. Alternatively, we could take the same approach as Fitzpatrick & Massa (1990) by adding additional parameters to the extinction function to infer our own dust relation. We could also extend the framework to provide population inferences on the surface gravity and effective temperature parameters. If we introduce Gaia parallaxes to our spectrophotometric framework, we could use the inferred surface gravity to estimate mass evolution of DA white dwarfs (Si et al., 2017, 2018). These ideas would not have been computationally feasible in the past, but are now within reach thanks to our GPU accelerated implementation of DAmodel.
An important science case that motivates the need for our all-sky accurate flux standards is SN Ia cosmology. Most recent efforts to constrain our Universe use compilations of SNe from a variety of surveys (Scolnic et al., 2022; Brout et al., 2022; Rubin et al., 2023). An essential part of the analysis is the cross-calibration of the surveys so they are on a consistent flux scale (Scolnic et al., 2015). Popovic et al. (2025) are already using our inferred SEDs in their Dovekie cross-calibration codebase. Their work will combine the present day’s largest SNe Ia samples from the Zwicky Transient Facility (Rigault et al., 2024), Dark Energy Survey (Abbott et al., 2024) and more, to further constrain the nature of our Universe.
6 Conclusions
We have developed a novel hierarchical Bayesian statistical framework for the spectrophotometric calibration of 32 faint DA white dwarfs ($19.516.5<V<$ 19.5) alongside three CALSPEC primary standards, from the UV to the NIR. We simultaneously utilise data from a variety of sources, including ground-based optical spectra, new space-based UV HST/STIS spectra and six HST/WFC3 photometric bands. In our framework we develop a flexible method of inferring a cubic spline surface to account for residual spectrograph instrumental systematics and inadequacies of the model. In parallel we jointly infer complex photometric systematics relating to time-varying sensitivity and count-rate nonlinearity. The model uses the most up to date data, DA WD templates and dust relations to allow for the most accurate modelling of the network yet; yielding an average HST/WFC3 RMS of less than 4 mmag from the UV to the NIR.
Including space-based HST/STIS UV spectra results in $A_{V}$ inferences that differ by an average of 10% compared to those derived without the UV data. We find that the additional UV coverage has little impact on $R_{V}$ inference, indicating that NIR spectra are also necessary for tighter constraints. Further improvements in modelling were achieved by adopting the dust relation from Gordon et al. (2023), which resulted in smaller residuals compared to the Fitzpatrick (1999) dust relation. By design, the hierarchical structure of our model does not affect individual DA WD parameter inferences, but instead enhances the precision of shared instrumental systematics estimation. When modelling HST/WFC3 photometry, we infer a NIR F160W count-rate nonlinearity of -3.19 $\pm$ 0.31 mmag/mag, in agreement with previous studies.
The model has seen vast performance improvements since Narayan et al. (2019) thanks to the combination of using Hamiltonian Monte Carlo, GPU acceleration and cubic splines rather than Gaussian Processes. This has resulted in convergence time going from a day per object to half an hour for the full sample. These speed-ups provide new opportunities to expand the model to include more WDs and further increase spectroscopic coverage. Additionally, we can apply the framework to larger samples for population-level parameter inference. The methods developed in this work will also inspire spectroscopic training of the next iteration of BayeSN, a hierarchical Bayesian model for SN Ia modelling (Mandel et al., 2009, 2011, 2022; Thorp et al., 2021; Grayling et al., 2024).
The faint all-sky DA white dwarf network will be imperative for the future of SN Ia cosmology. Work is already underway using the SEDs inferred from this work to combine today’s largest samples of SN Ia. With the addition of near-infrared spectra, the network will become even more useful in the calibration and synergies of current and next-generation surveys, including the James Webb Space Telescope, Vera Rubin Observatory’s Legacy Survey of Space and Time and the Nancy Grace Roman Telescope.
Acknowledgements
BMB is supported by the Cambridge Centre for Doctoral Training in Data-Intensive Science funded by the UK Science and Technology Facilities Council (STFC) and two G-Research early career researchers grants used for equipment and travel. BMB also acknowledges travel support provided by Christ’s College Cambridge’s Clayton Fund, LSST DESC supported by DOE funds administered by SLAC and STFC for UK participation in LSST through grant ST/S006206/1. BMB, AB, ML and GN were supported on this project through the HST Guest Observer Program (HST-GO-16764, PI G. Narayan). GN also gratefully acknowledges NSF support from AST-2206195, and a CAREER grant AST-2239364, supported in-part by funding from Charles Simonyi, and OAC-2311355, DOE support through the Department of Physics at the University of Illinois, Urbana-Champaign (#13771275), and HST-GO-17128 (PI R. Foley). Supernova and astrostatistics research at Cambridge University is supported by the European Union’s Horizon 2020 research and innovation programme under European Research Council Grant Agreement No 101002652 (PI K. Mandel) and Marie Skłodowska-Curie Grant Agreement No 873089.
Data Availability
The implementation of DAmodel is publically available at: https://github.com/benboyd97/DAmodel. All data used and our published SEDs can be found at: https://zenodo.org/records/14339961. The theoretical DA WD spectra (Hubeny et al., 2021) used to create the underlying grid can be found at: www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz. The SEDs used to tie our work to the CALSPEC system (Bohlin et al., 2020) may be found at: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/. The HST/WFC3 UVIS2 and IR filters (Calamida et al., 2022a) used in this work are available at: https://www.stsci.edu/hst/instrumentation/wfc3/performance/throughputs. Data from other surveys used to calculate residuals is published in Axelrod et al. (2023).
Appendix A Published SEDs
In Figure 12 we illustrate the SEDs for each WD in our network that we publish alongside our work. Each published WD SED is derived by averaging 2000 randomly drawn SEDs forward-modelled from 2000 posterior samples. With these SED draws we may also find the associated uncertainty at each wavelength found by taking the standard deviation at each pixel. Our published SEDs can be found at: https://zenodo.org/records/14339961.
Figure 12: Illustration of our 35 inferred SEDs for the DA WD network. The top three SEDs in blue are the three bright CALSPEC standards, while the rest in orange belong to the faint standards.
Appendix B Comparisons With Axelrod et al. (2023)
Axelrod et al. (2023) previously analysed the same network of stars. We plot the average ratio between our inferred SEDs and those from Axelrod et al. (2023) at wavelengths with mutual coverage in Figure 13. The largest differences are observed in the UV, likely driven by our inference of the $R_{V}$ parameter, the adoption of the updated dust relation and theoretical template, and the inclusion of HST/STIS UV spectra. Spikes in the optical are due to differences in inferred effective temperature, resulting in differing absorption feature strength.
In Figure 14 we compare parameter inferences between our model and Axelrod et al. (2023) when we use the same model assumptions. We see that when we adopt the Fitzpatrick (1999) dust relation, set a constant $R_{V}=3.1$ and ignore the HST/STIS UV spectra, the parameter inferences between the two models are consistent. This further supports the notion that differences in inferred SEDs are due to the differing model assumptions.
Figure 13: Ratios of SEDs inferred from this work and those from Axelrod et al. (2023). Differences come from using different theoretical SED templates, dust relations, $R_{V}$ assumptions, spectroscopic data and methodology. The blue line indicates the average ratio as a function of wavelength, while the grey shaded region shows the range of ratios across the 32 faint WDs.
Figure 14: Comparing parameter inferences when using the same modelling decisions as Axelrod et al. (2023). This included using optical spectra only, using the Fitzpatrick (1999) dust relation and a constant $R_{V}=3.1$.
Appendix C White Dwarf Parameter Correlations
In Figure 15, we present the posterior distributions of the object-level parameters for WDFS2317-29. The posterior shows strong positive correlations among the effective temperature $T_{\text{eff}}$, the dust extinction parameter $A_{V}$ and the achromatic offset parameter $\mu$. Higher extinction diminishes absorption features, leading the model to infer higher effective temperatures as compensation to match the observed data. This trade-off between effective temperature and extinction preserves the colour of the WD but alters its observed brightness, which is then subsequently adjusted through the positive correlation with the achromatic offset parameter $\mu$. Narayan et al. (2019) used a similar methodology and found the same three-dimensional posterior correlations in Figure 14 of their work. The Bayesian analysis in O’Callaghan et al. (2024) further explores trade-offs and degeneracies between dust and intrinsic stellar parameters using astrophysical priors and spectroscopic constraints.
Including the UV spectrum in the analysis raises the inferred effective temperature for WDFS2317-29, likely due to the additional information provided by the Lyman-$\alpha$ feature, which in turn increases the inferred extinction. As a result of the increased extinction and effective temperature, the inferred achromatic offset $\mu$ then rises to prevent an increase in the overall brightness of the synthetic photometry. The inclusion of STIS UV spectra, however, does not significantly affect the inference of $R_{V}$ for this object. Detecting differences in $R_{V}$ would require additional NIR spectra to capture the full panchromatic effects of dust.
Figure 15: Corner plot showing the inferred parameter posteriors for WDFS2317-29. Posteriors in red are inferred using only the optical spectra, while the blue posteriors use both optical spectra and STIS UV spectra. In the posteriors we see strong correlations between the effective temperature $T_{\text{eff}}$, the dust extinction parameter $A_{V}$ and the achromatic offset parameter $\mu$.
Appendix D HST/WFC3 Systematic Correlations
In Figure 16, we plot the inferred posteriors for the HST/WFC3 photometric zeropoints. Including the STIS UV spectra has no effect on the F475W band, because this is where we set the CALSPEC primary magnitudes to constant. Adding the STIS UV spectra encourages the F275W and F336W zeropoints to be lower and the F675W, F775W, and F160W zeropoints to be higher. The changes in zeropoints are a reflection of the average effect that including the STIS spectra has on the inferred SEDs, as illustrated in Figure 5.
We examine the correlation between dust and the inferred F160W CRNL in Figure 17. A positive correlation is observed between the average of inferred $A_{V}^{s}$ values and inferred CRNL value. This suggests that the CRNL inferred using this method is sensitive to the chosen dust relation and any additional assumptions applied to it. Similar sensitivity to the dust relation was previously noted in Section 4.3 when inferring a CRNL closer to the result reported in Axelrod et al. (2023) when using the Fitzpatrick (1999) dust relation.
Figure 16: HST/WFC3 Cycle 25 UVIS1 and IR zeropoint inferences using only optical spectra in red and STIS UV spectra (as well as optical) in blue.
Figure 17: Corner plot showing the average inferred $A_{V}^{s}$ and $R_{V}^{s}$ over 2000 posteriors samples along with the inferred HST/WFC3 F160W count-rate nonlinearity (CRNL). The plot shows a positive correlation between the average inferred $A_{V}^{s}$ and the inferred CRNL.
Appendix E Population Dust Constraints
Figure 18: Population dust parameters inferred with (blue) and without (red) the STIS UV spectra for the 19 WDs that have it. These results are only a demonstration of the hierarchical capabilities of the model and are influenced by selection effects. We show in black the constant population values motivated by literature that we use in the priors for our main analysis.
As a demonstration of our models ability to infer population-level parameters, we infer a $\mu_{RV}$ and $\sigma_{RV}$ given a truncated normal prior $R^{s}_{V}\sim TN(\mu_{RV},\sigma_{RV},1.2,\infty)$. We also infer an average population extinction $\tau$ assuming $A^{s}_{V}$ follows an exponential distribution. The resulting posteriors with and without using the HST/STIS UV spectra are shown in Figure 18. The inferred parameters are significantly different when compared to the constant population parameters used in our main analysis motivated by the work of Schlafly et al. (2016). The results in our population inference are to be taken with caution, as our DA WD network is strongly influenced by selection effects that are not taken into account in the analysis. In Section 5 we discuss future plans to apply the model to a larger volume limited sample of DA WDs to provide constraints on these population parameters that are uninfluenced by selection effects. The results show that the introduction of the HST/STIS UV spectra does not significantly change the $R_{V}$ population parameters. We would only expect substantial difference if we obtain NIR spectra for a panchromatic view to monitor the effects of the dust relations.
Appendix F HST Zeropoint Shifts
A key advantage of the hierarchical aspect of our work and the work of Axelrod et al. (2023) is that we can infer relative HST/WFC3 offsets that deviate from the absolute CALSPEC system. To derive the CALSPEC zeropoints for Cycle 25 (average MJD=57588.79 days) we adopt the same scheme as Narayan et al. (2019) where we use the difference between synthetic magnitudes and what is observed for the three primary stars only. In Table 4 we present zeropoints derived post-hoc by tying our relative zeropoints and those derived in Axelrod et al. (2023) to the F475W absolute zeropoint in the Bohlin (2014) and Bohlin et al. (2020) CALSPEC systems. Using the achromatic differences between the zeropoints in the table we can easily convert our network between systems without needing to reanalyse our data. In Figure 19 we plot the deviations from the two CALSPEC systems. We see that our work is in good agreement with results inferred by Axelrod et al. (2023) in the UV, with relative zeropoints deviating from CALSPEC zeropoints by around 11 mmag in the 2020 system and 4.5 mmag in the 2014 system. In the NIR we infer larger relative zeropoints shifts with CALSPEC with differences up to 7.5 mmag for the 2014 system and 15 mmag for the 2020 system. Despite an independent analysis, methodology, and dataset, our relative zeropoints show the same general trend as Axelrod et al. (2023), decreasing as a function of wavelength. We expect our relative zeropoints to differ of this order from CALSPEC since we have analysed a larger and fainter set of standards.
Figure 19: Relative HST/WFC3 zeropoint deviations away from the CALSPEC system for this work (presented as the solid blue line) and Axelrod et al. (2023) (presented as the orange dotted line). As a post-hoc step we arbitrarily to tie our relative zeropoints to the CALSPEC system at the F475W band using the difference between observed a synthetic magnitudes from Bohlin (2014) (top plot) and Bohlin et al. (2020) (bottom plot) for the three primary standards. The zeropoints calculated are AB magnitudes for UVIS1 and IR detectors during Cycle 25 (average MJD=57588.79 days) only.
Table 4: In this table we present inferred HST/WFC3 zeropoints for this work and Axelrod et al. (2023) in comparison to two CALSPEC systems.
CALSPEC Bohlin (2014)
F275W F336W F475W F625W F775W F160W
Boyd et al. $Z_{X}$ (mag) 23.9899 $\pm$ 0.0017 24.5553 $\pm$ 0.0012 25.5836 $\pm$ 0.0006 25.4182 $\pm$ 0.0009 24.7436 $\pm$ 0.0011 25.8004 $\pm$ 0.0016
Axelrod et al. $Z_{X}$ (mag) 23.9898 $\pm$ 0.0013 24.5596 $\pm$ 0.0008 25.5836 $\pm$ 0.0004 25.417 $\pm$ 0.0006 24.7481 $\pm$ 0.0008 25.8045 $\pm$ 0.0011
$Z_{X}^{\text{CALSPEC 14}}$ (mag) 23.9854 $\pm$ 0.0008 24.5565 $\pm$ 0.0007 25.5836 $\pm$ 0.0006 25.4173 $\pm$ 0.0007 24.7475 $\pm$ 0.0007 25.8078 $\pm$ 0.001
Boyd et al. $Z_{X}-Z_{X}^{\text{CALSPEC 14}}$ (mmag) 4.5 $\pm$ 1.8 -1.2 $\pm$ 1.4 0.0 $\pm$ 0.9 0.9 $\pm$ 1.2 -3.8 $\pm$ 1.3 -7.5 $\pm$ 1.9
Axelrod et al. $Z_{X}-Z_{X}^{\text{CALSPEC 14}}$ (mmag) 4.4 $\pm$ 1.5 3.1 $\pm$ 1.1 0.0 $\pm$ 0.8 -0.3 $\pm$ 0.9 0.6 $\pm$ 1.1 -3.4 $\pm$ 1.5
CALSPEC Bohlin et al. (2020)
F275W F336W F475W F625W F775W F160W
Boyd et al. $Z_{X}$ (mag) 23.9782 $\pm$ 0.0017 24.5437 $\pm$ 0.0012 25.5719 $\pm$ 0.0006 25.4066 $\pm$ 0.0009 24.732 $\pm$ 0.0011 25.7887 $\pm$ 0.0016
Axelrod et al. $Z_{X}$ (mag) 23.9784 $\pm$ 0.0013 24.5482 $\pm$ 0.0008 25.5722 $\pm$ 0.0004 25.4056 $\pm$ 0.0006 24.7367 $\pm$ 0.0008 25.7931 $\pm$ 0.0011
$Z_{X}^{\text{CALSPEC 20}}$ (mag) 23.9672 $\pm$ 0.0006 24.5395 $\pm$ 0.0007 25.5722 $\pm$ 0.0006 25.4089 $\pm$ 0.0007 24.7384 $\pm$ 0.0007 25.8041 $\pm$ 0.0009
Boyd et al. $Z_{X}-Z_{X}^{\text{CALSPEC 20}}$ (mmag) 11.0 $\pm$ 1.8 4.1 $\pm$ 1.4 -0.2 $\pm$ 0.9 -2.3 $\pm$ 1.2 -6.3 $\pm$ 1.3 -15.4 $\pm$ 1.8
Axelrod et al. $Z_{X}-Z_{X}^{\text{CALSPEC 20}}$ (mmag) 11.2 $\pm$ 1.4 8.6 $\pm$ 1.0 0.0 $\pm$ 0.7 -3.3 $\pm$ 0.9 -1.7 $\pm$ 1.1 -11.0 $\pm$ 1.4
Notes: We calculate two sets of zeropoints for both the Bohlin (2014) and the Bohlin et al. (2020) CALSPEC systems. CALSPEC zeropoints and uncertainties are inferred using the same scheme as Narayan et al. (2019) where we use the difference between synthetic magnitudes (derived from published SEDs) and what we observe in Cycle 25 for the three primary standards only. Our main results are already with respect to the CALSPEC 2020 system, tied to the F475W band. To move our results onto the CALSPEC 2014 system we simply add the difference between 2020 and 2014 systems in the F475W band to all bands. To place the Axelrod et al. (2023) relative zeropoints onto the CALSPEC 2014 system, we subtract their inferred changes in zeropoints (shifted so that the F475W is at zero) to the derived CALSPEC 2014 zeropoints. To then place Axelrod et al. (2023) zeropoints onto the CALSPEC 2020 system, we subtract the difference between 2020 and 2014 systems in the F475W band from all the bands. Presented zeropoints are equivalent to the UVIS1 and IR zeropoints during Cycle 25 (average MJD=57588.79 days) in AB magnitudes. Zeropoints are also calculated from aperture radii (7.5 pixels for UVIS and 5 pixels for the IR) to infinity.
Appendix G Inferred Spectroscopic Spline Residuals
To analyse the spectroscopic data of our network we add a cubic spline to the inferred SED in flux space. In Figure 20 we over-plot the inferred splines for both the UV STIS spectra and the optical spectra. We interpret the optical functions as reflecting instrumental systematics, with flux corrections being more biased near the spectrograph edges where throughput declines sharply. For the STIS UV spectra we see strong spline functions, particularly at shorter wavelengths. These results may be indicative of STIS systematics, resulting in an underestimation of flux. Alternatively, this may be understood as a difference between model and data, where perhaps our chosen dust relation needs improvement or our theoretical DA WD template needs refining. In Section 4.2 we discuss various STIS systematics and caveats in dust relations that may have led to these inferred splines.
Figure 20: Inferred spline residuals for the STIS UV spectra (left) and the ground-based optical spectra (right). Spline strength is expressed as the percentage of the maximum flux value in its respective spectrum. There is no meaning behind the colour of each spline.
Appendix H Residuals with different modelling assumptions
Table 5: HST/WFC3 residual mean $\mu$ and RMS for the 32 faint DA WDs when comparing their observed photometry with synthetic photometry for different experiments. Opt only signifies that we use only ground-based optical spectra and HST photometry to calibrate the stars. Opt+STIS signifies that we also use the STIS UV spectra for the 19 WDs that had coverage. G23 indicates that we use the Gordon et al. (2023) dust relation, whilst F99 represents the Fitzpatrick (1999) dust relation. Const $R_{V}$ refers to a constant $R_{V}=3.1$. Norm $R_{V}$ experiments used the truncated normal $R_{V}$ prior $TN(\mu_{RV}=3.1,\sigma_{RV}=0.18,1.2,\infty)$. Population $R_{V}$ experiments also inferred the population parameters $\mu_{RV}$, $\sigma_{RV}$ and $\tau$. We label tests presented in our final results with an *. Results shown are in AB millimagnitudes.
Test All F275W F336W F475W F625W F775W F160W
$\mu$ RMS $\mu$ RMS $\mu$ RMS $\mu$ RMS $\mu$ RMS $\mu$ RMS $\mu$ RMS
Opt Only G23 Const $R_{V}$0.01 4.36 -0.13 3.54 -0.11 3.44 1.52 4.34 -0.36 2.35 0.1 3.48 -1.09 7.3
Opt Only G23 Norm $R_{V}$0.04 3.79 0.02 2.25 -0.28 2.77 1.43 4.09 -0.37 2.3 0.22 3.44 -1.29 6.29
Opt Only G23 Pop $R_{V}$ 0.02 3.67 0.65 2.33 -0.81 2.65 1.12 3.74 -0.39 2.35 0.85 4.37 -1.29 5.47
Opt+STIS G23 Const $R_{V}$0.02 4.68 0.16 3.47 0.26 3.56 1.35 4.23 -0.64 2.4 0.03 3.44 -1.31 8.43
Opt+STIS G23 Norm $R_{V}$0.06 3.9 0.38 2.44 -0.07 3.29 1.2 3.78 -0.62 2.33 0.18 3.43 -1.44 6.56
Opt+STIS G23 Pop $R_{V}$0.07 3.6 1.16 3.03 -0.57 3.27 0.82 3.27 -0.67 2.4 0.42 3.38 -1.55 5.48
Opt Only F99 Const $R_{V}$ 0.05 4.59 1.42 4.27 -1.29 3.79 0.7 3.77 -0.21 2.32 0.56 3.51 -0.89 7.86
Opt Only F99 Norm $R_{V}$ 0.01 4.07 1.26 3.33 -1.17 3.13 0.64 3.77 -0.2 2.24 0.6 3.45 -1.09 6.9
Opt Only F99 Pop $R_{V}$0.01 3.85 1.3 3.39 -1.14 2.85 0.7 3.71 -0.18 2.2 0.5 3.33 -1.26 6.31
Opt+STIS F99 Const $R_{V}$0.06 4.81 1.38 4.35 0.04 3.8 -0.73 3.65 -0.67 2.5 1.23 3.61 -1.62 8.53
Opt+STIS F99 Norm $R_{V}$0.08 4.09 1.05 3.13 0.13 3.43 -0.66 3.5 -0.6 2.31 1.09 3.57 -1.47 6.96
Opt+STIS F99 Pop $R_{V}$0.08 3.84 0.84 3.19 0.33 3.08 -0.31 3.14 -0.43 2.25 0.7 3.24 -1.61 6.57
In our work we make various modelling assumptions that affect our ability to reproduce the data. An example of such a decision is whether we only use optical spectra in our inference, or whether we also use the STIS UV spectra for the 19 WDs that have coverage. Further assumptions were related to dust where we use the Gordon et al. (2023) relation, whereas previous analyses have used Fitzpatrick (1999). Finally, our priors on $R_{V}$ are important to consider, in previous analysis this has been kept constant as $R_{V}=3.1$, but in our work we use a truncated normal prior and also experiment with using a population prior. In Table 5 we show how varying these assumptions affects our residual bias and scatter with HST/WFC3 photometry.
We find that RMS scatter is lower overall when we use the updated Gordon et al. (2023) relation over the Fitzpatrick (1999) relation. Including the STIS UV spectra makes the RMS scatter marginally worse in all bands except the F475W. We also find that granting the model additional flexibility to infer population parameters for the $R_{V}$ prior significantly reduces the RMS scatter. Our goal is not solely to minimise RMS residuals with the HST/WFC3 photometry; rather, we aim to leverage all available data while adhering to physically motivated modelling principles. Accordingly, we present our final results based on analyses that incorporates both the optical and STIS UV spectra, utilises the updated Gordon et al. (2023) dust relation, and adopts a truncated normal $R_{V}$ prior informed by the work of Schlafly et al. (2016).
Appendix I Photometric Residuals with different surveys
Using our inferred synthetic SEDs for the WDs, we can integrate over various bandpass filters corresponding to different surveys. For any survey that observed a given standard, we can compute the residual between their measurement and our synthetic prediction. We show our residuals for DES in Figure 9, SDSS in Figure 21, PS1 in Figure 22, Gaia in Figure 23 and DECaLS in Figure 24. The data collection process and residual results are discussed in Section 4.4. Since each survey is calibrated independently, a wavelength-dependent bias is expected. Additional biases are anticipated because our SEDs are tied to the new CALSPEC Bohlin et al. (2020) system rather than Bohlin (2014). Quantifying these biases is crucial for cross-calibrating surveys and combining samples. In Section 5, we discuss the ongoing application of our SEDs to unify SN Ia samples across different surveys.
Figure 21: Photometric residuals when comparing our synthetic SDSS photometry with real SDSS observations (Holberg & Bergeron, 2006) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.
Figure 22: Photometric residuals when comparing our synthetic PS1 photometry with real PS1 observations (Flewelling et al., 2020) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.
Figure 23: Photometric residuals when comparing our synthetic Gaia photometry with real Gaia observations (Gaia Collaboration et al., 2023) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in Vega magnitudes.
Figure 24: Photometric residuals when comparing our synthetic DECaLS photometry with real DECaLS observations (Schlegel et al., 2021) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.